### v1

# Thermodynamic & algorithmic efficiency in heterogeneous networks

Abstract
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in a real world environment is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it. While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are gathered into a single quantity tantamount to the thermodynamic efficiency. We take as a case study an algorithm originally proposed for creating and searching trees of patterns under the assumption that, akin to living system, the algorithm seeks the best internal representation of the outer world, namely a real world dataset of patterns. Remarkably, optimal thermodynamic efficiency overlaps with the optimal algorithmic efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale in where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region. Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the critical-state when it confronts with higher-complexity environments, while it does not when the environment is simpler.
Keywords von Neumann entropy  $\cdot$ Specific heat  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning
1 Introduction
Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today. We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of optimal ability to store the information about the environment and the ability to transmit information across the structure [1, 2]. The latter is defined in terms of microscopic quantities thanks to the fairly newly developments in statistical physics of fields in network theory [2, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will than encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work a system seeks the best-fitness configuration given some external constraints $r$ and the dataset $\mathcal{D}$.
We take as case study the algorithm introduced by Portegys [11] for nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [12]. The algorithm employs a tree formation protocol where raw data are reorganized into tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing tree structure. At each step the new pattern is linked to either the currently parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the relative distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.  Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star-graph turns out. For some intermediate value of $r$ a ragged tree structure is produced, while for $r\rightarrow 2$ a 1d-lattice is created. Remarkably, the algorithm produces exclusively trees, i.e. number of spanning trees is always 1, which are sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection111 ”Natural selection does not work as an engineer works. It works like a tinkerer - who does not know exactly what he is going to produce but uses whatever he finds around him whether it be pieces of string, fragments of wood, or old cardboards; in short works like a tinkerer who uses everything at his disposal to produce some kind of workable object.”  [13], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a distance based rule.
Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
The Portegys algorithm has the desirable properties of being simple, explainable and subject to a single control parameter $r$. Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools as the von Neumann entropy and its derivative reflect the functional capabilities of the network. Most notably thermodynamic efficiency and algorithmic efficiency demonstrate noticeable overlap.
Figure 1: The three prototypical structures generated by the Portegys algorithm. The algorithm creates star-like networks up to 1d lattice when $r\rightarrow 2$. For intermediate values of $r$, networks resembling an arborescent tree are produced. The process is reminiscent of the transition vapor/liquid $\rightarrow$ amorphous solid $\rightarrow$ 1d lattice. Depicted networks are generated from a subset of 50 samples randomly extracted from the MNIST dataset with $r=0.05,0.95,1.50$ respectively. Classes are indicated by the 10 different colors in the color-bar.
2 Theoretical framework
2.1 Statistical physics of networks
Information spreading over networked structures is governed by the Laplacian matrix [14] and its spectrum of eigenvalues encodes many relevant topological properties of the graph [15, 16]. Let us consider an undirected and unweighted network $G(V,E)$ of order $|V|=N$ with number of links equal to $|E|=m$. The network Laplacian is defined as $\hat{L}=\hat{D}-\hat{A}$, where $\hat{D}$ and $\hat{A}$ are the diagonal matrix of the node degrees and the adjacency matrix respectively. The Laplacian operator can be used to define a density operator [5], where the Laplacian takes the role of the Hamiltonian of statistical physics [17] as in equation 1.
Given an initial state of the network encoding the concentration of quantity, e.g. information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$. Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counter-part of the path integral formulation for general diffusion processes [18, 19].
As $\hat{K}$ is a positive semidefinite Hermitian matrix admits the spectral decomposition in terms of the sum of projectors $\hat{K}=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}|$. Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$. This relation remarks that the eigenvectors of the Laplacian are akin to the wave-vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the ”wave length” $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3]. Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed defining a probability measure over the accessible modes decomposed by the spectral theorem. Such canonical ensemble [5, 4, 20] is encoded into the density matrix:
$\hat{\rho}_{\tau}=\frac{e^{-\tau\hat{L}}}{Z}$ (1)
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [18, 20]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a edge-centric random walker to be back in its initial location at time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
Finally, the network von Neumann entropy can be introduced [5] as a multi-scale descriptor of the network structure:
$S(\hat{\rho}(\tau))=-\sum_{i=1}^{N}\mu_{i}(\tau)\log\mu_{i}(\tau)$ (2)
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [20] as illustrated in figure 2. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the homogeneous phase where the information has propagated all over the network.
The network entropy thus counts the (logarithmic) number of informationally connected component on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [20, 3]. For instance, a constant specific heat reflects the scale-invariance of the network where information is lost at a constant rate. Such a situation is typical of d-dimensional lattices and scale invariant networks with power law distributed spectrum. On the contrary, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star-graph, depicted in figure 1, or an Erdös-Rényi network, see purple curve in figure 1. Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components is the full-network reviling a single trivial-scale.
From this considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 20]. In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction.
To conclude, we can intuitively use the diffusion parameter $\tau$ to resolve the network topological structures at different resolution scales. Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the ”wavelengths” of its modes.
Figure 2: Von Neumann entropy as a function of the diffusion time. Panel (a) depicts diffusion over two exemplary networks with analogous topological features to those depicted in panel (b): in blue a stochastic block model network exhibiting tripartite modular structure, and in purple a Erdös-Rényi random network with the same number of nodes of the blue one. For small $\tau\sim 0$, both networks exhibit the maximum spectral entropy value $\log N$, where $N$ is the number of nodes, reflecting the segregated phase. For $\tau\rightarrow\infty$, the spectral entropy converges to $\log C$, where $C$ is the number of connected components, thus reflecting the integrated phase. For intermediate values of $\tau$, the Erdös-Rényi network shows a single critical scale at which the network transitions from the segregated to the integrated phase. The position on the horizontal axis of the critical scale is related to the edge density of the network as well as the heterogeneity in the degree distribution. On the contrary, the blue network exhibits two critical resolution scales in the entropic transition. At smaller $\tau$, information initially diffuses within the modules ending into a plateau. A later slower diffusion process between the three modules completes the transition to the integrated phase. The insets show the density matrix $\hat{\rho}_{\tau}$ across the entropic transition on the blue curve. The tripartite modular structure of the network is shown by the emergence of the three blocks in the density matrix. Graphs depicted in (b) have only illustration purpose. Illustration from [21].
2.2 Thermodynamics of networks
The beauty of the statistical physics theory for networks introduced above is that it naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities as work, $W$, and heat, $Q$, can be defined in (at least mathematical) analogy to their thermodynamic counterparts [2]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network. In turns, lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
$Q=-\frac{dS}{\tau}=-\frac{\log N-S}{\tau}\leq 0$ (3)
which resembles the heat dissipation in a thermodynamic process.
On the other hand, it is obvious that adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [2, 4]. Less obvious is how, for fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [14, 22]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq\log N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
$W=dF=F-F_{iso}=F+\frac{\log N}{\tau}\geq 0$ (4)
which is tantamount to the work in the canonical ensemble. Moreover, $U=W+Q\geq 0$ which implies $W\geq|Q|$. A formation process can be depicted as the trade-off between this two necessities of a networked complex systems: the maximization of information flow, while at the same time maintening an amount of entropy such that the network can encode and express sufficient amount of diversity trough its traits.
Thus, to complete the analogy to thermodynamics, the trade-off between diversity of information and information flow is captured by the efficiency [2]:
$\eta=\frac{W+Q}{W}=1-\frac{|Q|}{W}=\frac{U}{W}$ (5)
where $0\leq\eta\leq 1$ and $U\equiv-\partial_{\tau}\log Z\equiv\operatorname{Tr}[\hat{L}\hat{\rho}]$. The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness and heterogeneity [2]. In our work we make use of this variational principle to explain the structural properties of a network that is required to encode the diversity expressed in the environment, where the environment is defined by a real-world dataset of patterns. We also make use of the specific heat to identify which is the resolution scale at which such variational principle is effective for heterogeneous networks constrained to a sparsity regime.
3 Results
Our results are evaluated over subsets of three different datasets: the NIST [23], the MNIST [24] and the FashionMNIST [25] datasets. These data sets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure. The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep- learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently. In the Suppplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashioMNIST. The 10 classes are equally represented both in the test and in training set. We limit to $r\leq 0.85$ as the accuracy, and hence the algoritmic efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
3.1 Algorithmic efficiency
After the tree formation described in section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the performance would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure. The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$. The distance is then compared to the distance between the test pattern and all of the root node’s children. The first node with a pattern that has a smaller distance is than selected as the new root node, and the process starts over with the newly selected node as root. The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
It is natural to define an algorithmic efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
$\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
the logarithm is taken as the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes. For the case of a star-graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of figure 3.
Figure 3: Algorithmic efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best-match, (b). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency. The shaded grey area designates the standard deviation. The purple shaded area in panel (c) indicates the intersection between the blue curve and the standard deviation at the optimal point whose y-axis projection is indicated by yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset. The 10 digit-classes are equally represented both in the train- and test-set.
3.2 Thermodynamic and algorithmic efficiency at the emergence of new C-peaks
In figures 4, 6 and 5 methodology and results showing the coherence between algorithmic and thermodynamic efficiency are replicated for the three investigated datasets. For each dataset, panel (a) depicts the $\tau$-resolution scales identified by the first peak (blue curve), $\tau^{(1)}_{peak}$, and last peak (green curve), $\tau^{(2)}_{peak}$, of the specific heat, $C$, as a function over the control parameter $r$. See figure S5 in the Supplementary for more von Neumann entropy and specific heat curves as a function of $r$.
Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star-graph. Thus, structure emerges reflected by the arise of a second peak in $C$ and a roughly $C=c_{0}>0$ plateau in between the first and the last peak. Such transition can be identified by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ respectively.
Before the bifurcation the two curves overlap, and the single peak in the specific heat identifies a single resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure. For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice [3]. The observation is cross-validated by the color code in panel (d) showing the specific heat as a function of both parameters $r$ and $\tau$. The two curves $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ are projected onto the specific heat surface in panel (d).
In panels (b), (c) the overlap between the thermodynamic efficiency, $\eta$, and and the algorithmic efficiency, $\theta$, is illustrated. The dashed lines indicate where best trade-off is obtained, while the shaded purple region indicate a tolerance region for $\theta$ defined by the intersection between the points on the efficiency curve and the standard deviation computed at the best trade-off. Striking overlap between the two efficiency is observed for the NIST and the FashionMNIST dataset, suggesting that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles. Remarkably, the efficiency $\eta$ is computed at the $\tau$-resolution scales in the proximity of the transition depicted in panels (a), namely at the $\tau$ values where the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ appears in panel (a).
In the seminal work [2], it was indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this might give the time-scale where the diffusion process has reached equilibrium and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales. Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a) for all three figures 4, 6 and 5 show the numerical average $<\tau_{diff}>$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above. While the mentioned strategy is effective in the case of a transition from sparse to connected networks [2], it is less so in a transition that exclusively involves the heterogeneity of the degree distribution. We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta((N-1)/|E|-1)$.
Thus, it is crucial to identify which is the $\tau$-scale involved in the phase transition from a single trivial scale, i.e. the star-graph, to more complex topological traits. Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In section S3 of the Supplementary the efficiency, $\eta$, is computed in the range interval indicated by $<\tau_{diff}>$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
Figure 4: Thermodynamic & algorithmic efficiency - NIST. In (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the transition from the single-peak to $C=c_{0}>0$ plateau reflecting the emergence of a roughly scale-invariant structure. The numerical average of the diffusion time $\tau_{diff}=<1/\lambda_{2}>$ is plotted in red for reference, see the main text for discussion. In (d) the specific heat, $C$, surface is illustrated and the $\tau^{(1,2)}_{peak}$ are projected on it. In (b)-(c) the thermodynamic efficiency, $\eta$, and the algorithmic efficiency, $\theta$, are functions of $r$. Vertical dashed-lines indicate the global maxima of the efficiencies. Remarkably, the thermodynamic efficiency is computed in the $\tau$-range of one order of magnitude from the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ in (a). Note that both maxima in the efficiencies are located where a discontinuity in $\tau^{(2)}_{peak}$ is observed at $r=0.7$, far away from the transition. This is in line with the argument that indicates the NIST as a fairly simple dataset that does not benefit from the critical point. Curves are averaged over 50 independent simulations.
Figure 5: Thermodynamic & algorithmic efficiency - FashionMNIST. As for the case of the NIST, in (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the topological transition. In red the numerical average of the diffusion time $<\lambda_{2}=1/\tau_{diff}>$ is plotted for reference, see the main text for discussion. In (d) the specific heat, $C$, surface is illustrated. In (b)-(c) the efficiencies $\eta$ and $\theta$ are plotted as a function of $r$ with the dashed-lines indicating the global maxima. The $\eta$ is computed in the $\tau$-range of one order of magnitude from the $\tau$-scale indicated by the bifurcation of $\tau^{(1,2)}_{peak}$ in (a). Note that the maxima of the efficiencies is posited at the edge of the critical value of $r\sim 0.35$ indicating the FashionMNIST as a dataset more complex than the other two investigated, that does benefit from the critical region. Curves are averaged over 50 independent simulations.
Figure 6: Thermodynamic & algorithmic efficiency - MNIST. See the main text and the caption of figure 4 for a detailed explanation. The MNIST does not fully benefit from the transition point $r=0.3$, nevertheless we still observe a local maxima in the efficiency. The contraction of the $\tau^{(2)}_{peak}$ is reflected in the local maxima of $\eta$ in (b), but not in the efficiency (c). Overlap between the two efficiencies is less striking when compared to figures 4 and 5 but still valid if we include points in one standard deviation, see shaded purple area in (c), from the maximum of $\theta$. Note that the shaded area in (c) extends and includes local to global maxima in (c). Curves are averaged over 50 independent simulations.
3.3 Criticality or not to criticality
We now focus the discussion of the results obtained for the NIST and MNIST datasets. As it is shown in both panels (a) and (b) of figures 4 and 6, the ”plateau” region of the specific heat has a contraction and an expansion for $r=0.7$ for the NIST and $r=0.6$ for the MNIST. For the latter the contraction is such that the two $\tau^{(1,2)}_{peak}$ curves overlap. We mention that this effect is mitigated for larger datasets, i.e. larger networks, as we show in section S2 of the Supplementary. Moreover, for the MNIST case, the contraction of the plateau region is sufficiently strong to reflect on a local maximum in the efficiency, $\eta$, but not registered by the algorithmic efficiency $\theta$.
Note that the mentioned values of $r$ where the (largest) jump in the $\tau^{(2)}_{peak}$ is observed correspond to the $r$ values where both the best algorithmic efficiencies, $\eta$ and $\theta$, are found, as indicated by the red dashed lines in panels (c) and the multi-colored dashed lines in panels (b) (valid also for the MNIST if we include points in the one standard deviation illustrated by the shaded-purple). Nevertheless, we stress that we did not compute the efficiency at the corresponding $\tau$ scale where such discontinuity appears.
[…]
Moreover, while the value of $r$ where a structural transition appears is dependent on the dataset, the $\tau\sim 10^{1}$ scale identified is the same for all the three datasets. Interestingly that $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary S3). We speculate about the possibility that this is due to the constraint on sparsity.
4 Discussion
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [27]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset- they are posed in contact with. Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or in the very network-features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [28], or in a mixed-way as in artificial neural networks where the engineer designs a complicated distribution through the definition of an architecture and the training phase fills in the parameters in order to obtain the desired function-approximation machine to achieve statistical generalization [29].
When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [30]. Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to a task can be reported, e.g. class-matching. By using the recently developed theoretical tools of statistical physics for networks [5, 4, 31, 20, 3, 2], we have quantified diversity in the traits of the network with the von Neumann entropy and transmission flow of information as the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities [2], the formation process of a networked structure is depicted as the trade-off between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other. We have shown that such a thermodynamic efficiency strikingly overlaps with the algorithmic efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$. We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution involved in the transition from the trivial-structure with a single specific heat peak to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate. Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the hetrogeneity-transition from star-graph to 1d-lattice. While the transition point reflects the network resolution scale where the efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency. We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower. This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [26, 10].
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing,. LS: Writing – review & editing, Conceptualization, Supervision, Funding acquisition.
Funding
This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program ”Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No. 861153, as well as by the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).
Acknowledgments
We are thankful to Jos van Goor for providing some of the scripts that enabled efficient grid-search across parameters and to Federico Balducci for the useful discussions.
Supplemental Data
Supplementary Material include figures that are mentioned in the main text but are omitted from it to the purpose of clarity.
Data Availability Statement
The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s. Datasets and scripts can be found at https://github.com/CipolliniDavide/LaplacianTrees.

### v2

@@
-Thermodynamic & algorithmic efficiency in heterogeneous networks
+Thermodynamic & pattern-matching efficiency in heterogeneous networks
 Abstract
 Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today.
-At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in a real world environment is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it.
-While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are gathered into a single quantity tantamount to the thermodynamic efficiency.
+At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in the outer world is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it.
+While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are combined into a single quantity tantamount to the thermodynamic efficiency.
 We take as a case study an algorithm originally proposed for creating and searching trees of patterns under the assumption that, akin to living system, the algorithm seeks the best internal representation of the outer world, namely a real world dataset of patterns.
-Remarkably, optimal thermodynamic efficiency overlaps with the optimal algorithmic efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale in where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region.
-Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the critical-state when it confronts with higher-complexity environments, while it does not when the environment is simpler.
+Remarkably, the optimal thermodynamic efficiency overlaps with the optimal pattern-matching efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region.
+Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the vicinity to the critical point whether it is faced with higher-complexity environments, while it does not when the environment is simpler.
 Keywords von Neumann entropy  $\cdot$ Specific heat  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning
@@
 Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today.
-We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of optimal ability to store the information about the environment and the ability to transmit information across the structure [1, 2].
+We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [1, 2].
 The latter is defined in terms of microscopic quantities thanks to the fairly newly developments in statistical physics of fields in network theory [2, 3, 4, 5].
@@
 How these traits are optimized to maximize the fitness to the environment is an open question.
-The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines.
+The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines.
 Here we challenge this hypothesis.
-In our work a system seeks the best-fitness configuration given some external constraints $r$ and the dataset $\mathcal{D}$.
-We take as case study the algorithm introduced by Portegys [11] for nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [12].
+In our work a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
+We take as case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [13].
 The algorithm employs a tree formation protocol where raw data are reorganized into tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing tree structure.
@@
 Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$.
-For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star-graph turns out.
+For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out.
 For some intermediate value of $r$ a ragged tree structure is produced, while for $r\rightarrow 2$ a 1d-lattice is created.
@@
 The Portegys algorithm has the desirable properties of being simple, explainable and subject to a single control parameter $r$.
-Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools as the von Neumann entropy and its derivative reflect the functional capabilities of the network.
-Most notably thermodynamic efficiency and algorithmic efficiency demonstrate noticeable overlap.
+Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network.
+Most notably thermodynamic efficiency and pattern-matching efficiency demonstrate noticeable overlap.
 Figure 1: The three prototypical structures generated by the Portegys algorithm.
@@
 information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$.
-Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counter-part of the path integral formulation for general diffusion processes [18, 19].
+Where $\hat{K}=e^{-\tau\hat{L}}$ is the network propagator and is the discrete counterpart of the path integral formulation for general diffusion processes [19, 20].
 As $\hat{K}$ is a positive semidefinite Hermitian matrix admits the spectral decomposition in terms of the sum of projectors $\hat{K}=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}|$.
@@
 Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [19, 21].
-The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a edge-centric random walker to be back in its initial location at time $\tau$ [4].
+The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location at time $\tau$ [4].
 Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
 Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the homogeneous phase where the information has propagated all over the network.
-The network entropy thus counts the (logarithmic) number of informationally connected component on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [20, 3].
+The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [21, 3].
 For instance, a constant specific heat reflects the scale-invariance of the network where information is lost at a constant rate.
@@
 On the contrary, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network.
-The latter is the typical situation of a star-graph, depicted in figure 1, or an Erdös-Rényi network, see purple curve in figure 1.
-Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components is the full-network reviling a single trivial-scale.
+The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network, see purple curve in Figure 1.
+Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components components are the full-network revealing a single trivial-scale.
 From this considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 21].
@@
 Graphs depicted in (b) have only illustration purpose.
-Illustration from [21].
+Illustration from [22, 21, 23].
 2.2 Thermodynamics of networks
-The beauty of the statistical physics theory for networks introduced above is that it naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables.
+The statistical physics theory for networks introduced above naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables.
 To this regard network thermodynamic quantities as work, $W$, and heat, $Q$, can be defined in (at least mathematical) analogy to their thermodynamic counterparts [2].
@@
 On the other hand, it is obvious that adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [2, 4].
-Less obvious is how, for fixed number of edges, the degree distribution influences the flow of information.
+Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information.
 As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [15, 24].
 As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter.
-From $1\leq Z\leq\log N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$.
+From $1\leq Z\leq N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$.
 Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
@@
 Moreover, $U=W+Q\geq 0$ which implies $W\geq|Q|$.
-A formation process can be depicted as the trade-off between this two necessities of a networked complex systems: the maximization of information flow, while at the same time maintening an amount of entropy such that the network can encode and express sufficient amount of diversity trough its traits.
+A formation process can be depicted as the trade-off between these two necessities of the networked complex system: the need to maximize the information flow, while at the same time retaining an amount of structural entropy such that the network can encode and express sufficient diversity through its traits.
 Thus, to complete the analogy to thermodynamics, the trade-off between diversity of information and information flow is captured by the efficiency [2]:
@@
 We also make use of the specific heat to identify which is the resolution scale at which such variational principle is effective for heterogeneous networks constrained to a sparsity regime.
-3 Results
+3 Methods
 Our results are evaluated over subsets of three different datasets: the NIST [25], the MNIST [26] and the FashionMNIST [27] datasets.
@@
 In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently.
-In the Suppplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashioMNIST.
+In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST.
 The 10 classes are equally represented both in the test and in training set.
-We limit to $r\leq 0.85$ as the accuracy, and hence the algoritmic efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
-3.1 Algorithmic efficiency
+We limit to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
+The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] and where distance between patterns is the Euclidean distance.
+3.1 Pattern-matching efficiency
 After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach.
-The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the performance would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure.
+The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure.
 The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$.
@@
 The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
-It is natural to define an algorithmic efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
+It is natural to define an pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
 $\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
-the logarithm is taken as the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes.
-For the case of a star-graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of figure 3.
-Figure 3: Algorithmic efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best-match, (b).
+the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes.
+For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of Figure 3.
+Figure 3: Pattern-matching efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best match, (b).
 The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency.
 The shaded grey area designates the standard deviation.
-The purple shaded area in panel (c) indicates the intersection between the blue curve and the standard deviation at the optimal point whose y-axis projection is indicated by yellow shaded area.
+The purple shaded area in panel (c) indicates a tolerance $r$-domain of optimal efficiency.
+The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by yellow shaded area.
 Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset.
 The 10 digit-classes are equally represented both in the train- and test-set.
-3.2 Thermodynamic and algorithmic efficiency at the emergence of new C-peaks
-In figures 4, 6 and 5 methodology and results showing the coherence between algorithmic and thermodynamic efficiency are replicated for the three investigated datasets.
-For each dataset, panel (a) depicts the $\tau$-resolution scales identified by the first peak (blue curve), $\tau^{(1)}_{peak}$, and last peak (green curve), $\tau^{(2)}_{peak}$, of the specific heat, $C$, as a function over the control parameter $r$.
-See figure S5 in the Supplementary for more von Neumann entropy and specific heat curves as a function of $r$.
-Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star-graph.
-Thus, structure emerges reflected by the arise of a second peak in $C$ and a roughly $C=c_{0}>0$ plateau in between the first and the last peak.
-Such transition can be identified by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ respectively.
-Before the bifurcation the two curves overlap, and the single peak in the specific heat identifies a single resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph.
+4 Results
+4.1 Thermodynamic and pattern-matching efficiency at the emergence of new C-peaks
+Figure 4 shows the coherence between algorithmic defined in Section 3 and thermodynamic efficiency in Equation 5 for the three investigated datasets.
+In panels (a)-(c) the pattern-matching efficiency, $\theta$, is depicted.
+The vertical red-dashed line indicates the $r$ value for the optimal algorithmic trade-off, while the shaded purple region indicates a $r$-domain tolerance of the optimal pattern-matching efficiency.
+The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
+In panels (d)-(f) the thermodynamic efficiency, $\eta$, shows remarkable overlap with the pattern-matching efficiency $\theta$ in terms of optimal $r$.
+See the $r$-range spanned by the purple-shaded areas in panels (a)-(c) and the vertical dashed lines in panels (d)-(f).
+The overlap between the two efficiencies suggests that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles.
+See the next Section for the discussion on the double peak of the $\eta$ curves for the MNIST.
+Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph.
+Thus, the emergence of structure is reflected by the arise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (e)-(i).
+Such transition is also depicted in Figure 5 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (g)-(i) of Figure 4.
+Before the bifurcation the two curves overlap and the single peak in the specific heat identifies a single trivial resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph.
 The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure.
-For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice [3].
-The observation is cross-validated by the color code in panel (d) showing the specific heat as a function of both parameters $r$ and $\tau$.
-The two curves $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ are projected onto the specific heat surface in panel (d).
-In panels (b), (c) the overlap between the thermodynamic efficiency, $\eta$, and and the algorithmic efficiency, $\theta$, is illustrated.
-The dashed lines indicate where best trade-off is obtained, while the shaded purple region indicate a tolerance region for $\theta$ defined by the intersection between the points on the efficiency curve and the standard deviation computed at the best trade-off.
-Striking overlap between the two efficiency is observed for the NIST and the FashionMNIST dataset, suggesting that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles.
-Remarkably, the efficiency $\eta$ is computed at the $\tau$-resolution scales in the proximity of the transition depicted in panels (a), namely at the $\tau$ values where the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ appears in panel (a).
-In the seminal work [2], it was indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this might give the time-scale where the diffusion process has reached equilibrium and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
+For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice passing by $c_{0}=4/6$ for random trees notoriously related to the spectral dimension $d=2c_{0}$ [3].
+See also Section S4 illustrating the spectral entropy and specific heat curves for the three datasets independently depicted for each $r$-value.
+Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the structural transition.
+Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5.
+In the seminal work [2], it is indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
+the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium according to the authors, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
 Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$.
-Panels (a) for all three figures 4, 6 and 5 show the numerical average $<\tau_{diff}>$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above.
-While the mentioned strategy is effective in the case of a transition from sparse to connected networks [2], it is less so in a transition that exclusively involves the heterogeneity of the degree distribution.
-We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta((N-1)/|E|-1)$.
-Thus, it is crucial to identify which is the $\tau$-scale involved in the phase transition from a single trivial scale, i.e.
-the star-graph, to more complex topological traits.
+Panels (a)-(c) of Figure 5 show the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above.
+The key difference is in the use of the ensemble average which employs all the eigenvalues to compute the specific heat.
+Thus, it is crucial to identify which is the $\tau$-scale which resolves the phase transition from the single trivial scale of the star graph to more complex topological traits of the amorphous network depicted in Figure 1.
 Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
-In section S3 of the Supplementary the efficiency, $\eta$, is computed in the range interval indicated by $<\tau_{diff}>$, i.e.
+In Section S3 the efficiency, $\eta$, is computed in the range interval indicated by the diffusion time as defined in [2] $\tau_{diff}=1/\lambda_{2}$, i.e.
 the red curves depicted in panels (a).
 Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
-Figure 4: Thermodynamic & algorithmic efficiency - NIST.
-In (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the transition from the single-peak to $C=c_{0}>0$ plateau reflecting the emergence of a roughly scale-invariant structure.
-The numerical average of the diffusion time $\tau_{diff}=<1/\lambda_{2}>$ is plotted in red for reference, see the main text for discussion.
-In (d) the specific heat, $C$, surface is illustrated and the $\tau^{(1,2)}_{peak}$ are projected on it.
-In (b)-(c) the thermodynamic efficiency, $\eta$, and the algorithmic efficiency, $\theta$, are functions of $r$.
+Moreover note in Figure 5 the slow down of diffusion indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition.
+We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
+Figure 4: Thermodynamic & pattern-matching efficiency.
+In (a)-(c) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the three investigated datasets of increasing complexity (see arrow at the bottom of the figure).
+The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line.
+In (d)-(f) the optimal thermodynamic efficiency, $\eta$, as a function of $r$ overlaps with the pattern-matching efficiency.
 Vertical dashed-lines indicate the global maxima of the efficiencies.
-Remarkably, the thermodynamic efficiency is computed in the $\tau$-range of one order of magnitude from the bifurcation between $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ in (a).
-Note that both maxima in the efficiencies are located where a discontinuity in $\tau^{(2)}_{peak}$ is observed at $r=0.7$, far away from the transition.
-This is in line with the argument that indicates the NIST as a fairly simple dataset that does not benefit from the critical point.
-Curves are averaged over 50 independent simulations.
-Figure 5: Thermodynamic & algorithmic efficiency - FashionMNIST.
-As for the case of the NIST, in (a) the bifurcation of $\tau^{(2)}_{peak}$ and $\tau^{(2)}_{peak}$ curves identify the $\tau$-scale relevant for the topological transition.
-In red the numerical average of the diffusion time $<\lambda_{2}=1/\tau_{diff}>$ is plotted for reference, see the main text for discussion.
-In (d) the specific heat, $C$, surface is illustrated.
-In (b)-(c) the efficiencies $\eta$ and $\theta$ are plotted as a function of $r$ with the dashed-lines indicating the global maxima.
-The $\eta$ is computed in the $\tau$-range of one order of magnitude from the $\tau$-scale indicated by the bifurcation of $\tau^{(1,2)}_{peak}$ in (a).
-Note that the maxima of the efficiencies is posited at the edge of the critical value of $r\sim 0.35$ indicating the FashionMNIST as a dataset more complex than the other two investigated, that does benefit from the critical region.
-Curves are averaged over 50 independent simulations.
-Figure 6: Thermodynamic & algorithmic efficiency - MNIST.
-See the main text and the caption of figure 4 for a detailed explanation.
-The MNIST does not fully benefit from the transition point $r=0.3$, nevertheless we still observe a local maxima in the efficiency.
-The contraction of the $\tau^{(2)}_{peak}$ is reflected in the local maxima of $\eta$ in (b), but not in the efficiency (c).
-Overlap between the two efficiencies is less striking when compared to figures 4 and 5 but still valid if we include points in one standard deviation, see shaded purple area in (c), from the maximum of $\theta$.
-Note that the shaded area in (c) extends and includes local to global maxima in (c).
-Curves are averaged over 50 independent simulations.
-3.3 Criticality or not to criticality
-We now focus the discussion of the results obtained for the NIST and MNIST datasets.
-As it is shown in both panels (a) and (b) of figures 4 and 6, the ”plateau” region of the specific heat has a contraction and an expansion for $r=0.7$ for the NIST and $r=0.6$ for the MNIST.
-For the latter the contraction is such that the two $\tau^{(1,2)}_{peak}$ curves overlap.
+Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (g)-(i).
+The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure.
+Note that for datasets of higher complexity the highest efficiency, both thermodynamic and algorithmic, are posited at the critical value of the structural phase transition illustrated by the specific heat in panels (g)-(i).
+This is in line with the argument that states that a fairly simple dataset does not benefit from the critical point.
+Quantities are averaged over 50 independent simulations.
+Figure 5: Role of the ensemble average on the estimation of the $\tau$-scale.
+Panels (a)-(c) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$.
+$\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 4.
+$\tau_{diff}$ is the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$.
+Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (g)-(i) of Figure 4.
+At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$.
+The equilibrium of the diffusion process is an increasing function of $r$.
+$\tau_{diff}$ curves are averaged over 50 independent simulations.
+The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
+4.2 Performance at criticality
+As previously discussed, the $\tau$-scale where the bifurcation in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5 shows up indicates the scale involved in the structural transition and it is the resolution scale where the thermodynamic efficiency overlaps with the pattern-matching efficiency as a function of $r$.
+Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
+To discuss this empirical result we turn our attention to the panels (a)-(c) in Figure 4 and panels (a)-(c) in Figure 5.
+While for the FashionMNIST dataset there is a rough correspondence between the $r\sim 0.30$ where the structural transition emerges and the $r$-range $r\in(0.30,0.50)$ where the global maximum of the efficiencies is observed, for the MNIST and NIST datasets the thermodynamic and algorithmic efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.35$ and $r\sim 0.45$.
+The term ”far away” is quantified by the purple-shaded areas Figure 4.
+We rather observe in panels (b), (c), (e) and (f) of Figure 4 that the maximum efficiencies are posited where the $\tau^{(2)}_{peak}$ have a discontinuous jump in panels (b), (c) of Figure 5 at $r\sim 0.60$ for the MNIST and at $r\sim 0.70$ for the NIST.
+The argument holds also for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
+The jump for the MNIST curve in Figure 5 is such that the two $\tau^{(1,2)}_{peak}$ curves overlap.
 We mention that this effect is mitigated for larger datasets, i.e.
 larger networks, as we show in Section S2 of the Supplementary.
-Moreover, for the MNIST case, the contraction of the plateau region is sufficiently strong to reflect on a local maximum in the efficiency, $\eta$, but not registered by the algorithmic efficiency $\theta$.
-Note that the mentioned values of $r$ where the (largest) jump in the $\tau^{(2)}_{peak}$ is observed correspond to the $r$ values where both the best algorithmic efficiencies, $\eta$ and $\theta$, are found, as indicated by the red dashed lines in panels (c) and the multi-colored dashed lines in panels (b) (valid also for the MNIST if we include points in the one standard deviation illustrated by the shaded-purple).
-Nevertheless, we stress that we did not compute the efficiency at the corresponding $\tau$ scale where such discontinuity appears.
+The jump is caused by the presence of a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Figure 4 and the entropy curves depicted in Figure S6).
+Moreover, for the MNIST case, the jump is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$.
+Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 4(b), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 4(e).
+Nevertheless, we stress once again that we did not compute the thermodynamic efficiency at the corresponding $\tau$ scale where discontinuity of the $\tau^{(2)}_{peak}$ curve appears, but we rather measured $\eta$ at the $\tau$ scale indicated by the bifurcation.
+To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or father away, we turn to the argument, already present in the literature [28, 10, 29], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only weather the task requires a high level of complexity.
+While for the FashionMNIST, a notoriously more complex dataset when compared to both the MNIST and NIST, the optimal value for the efficiency is posited in the proximity of the critical value of $r\sim 0.30$, the optimal values of $r$ for the NIST and MNIST datasets are evidently far away from their respective transition values.
+The MNIST, which is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$.
+While the global maximum in Figure 4(e) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 5, the local maxima is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation.
+This observation is in line with the hypothesis that the system takes advantage from the critical behaviour only when a fairly complex task is presented to it and in line with other observations discussed in the literature [28, 10].
+Thus, it stresses the importance of dataset under analyses.
 […]
-Moreover, while the value of $r$ where a structural transition appears is dependent on the dataset, the $\tau\sim 10^{1}$ scale identified is the same for all the three datasets.
-Interestingly that $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary S3).
-We speculate about the possibility that this is due to the constraint on sparsity.
 5 Discussion
@@
 Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources.
-In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to a task can be reported, e.g.
+In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g.
 class-matching.
-By using the recently developed theoretical tools of statistical physics for networks [5, 4, 31, 20, 3, 2], we have quantified diversity in the traits of the network with the von Neumann entropy and transmission flow of information as the Helmholtz free energy.
-As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities [2], the formation process of a networked structure is depicted as the trade-off between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other.
-We have shown that such a thermodynamic efficiency strikingly overlaps with the algorithmic efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
+By using the recently developed theoretical tools of statistical physics of fields for networks [5, 4, 34, 21, 3, 2], we have quantified the diversity in the traits of the network with the von Neumann entropy and the transmission flow of information with the Helmholtz free energy.
+As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities, the formation process of a complex networked structure is interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [2].
+We have shown that such a thermodynamic efficiency strikingly overlaps with the pattern-matching efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
 Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$.
-We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution involved in the transition from the trivial-structure with a single specific heat peak to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
-Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the hetrogeneity-transition from star-graph to 1d-lattice.
-While the transition point reflects the network resolution scale where the efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency.
+We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial-structure, that is a single specific heat peak, to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
+Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the heterogeneity-transition from star graph to 1d-lattice trough amorphous.
+While the transition point reflects the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency.
 We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower.
-This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [26, 10].
+This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [28, 10, 29].
 Conflict of Interest Statement
@@
 Author Contributions
-DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing,.
+DC: Theoretical framework, Conceptualization, Methodology, Simulations, Data curation, Software, Writing – original draft, Writing – review & editing.
 LS: Writing – review & editing, Conceptualization, Supervision, Funding acquisition.
@@
 The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s.
-Datasets and scripts can be found at https://github.com/CipolliniDavide/LaplacianTrees.
+Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.

### v3

@@
-Thermodynamic & pattern-matching efficiency in heterogeneous networks
+Tree networks of real-world data: analysis of efficiency and spatiotemporal scales
 Abstract
-Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today.
-At the border between statistical physics, machine learning and network theory, we provide some computational empirical evidence indicating that the emergence of complex structures in the outer world is the macroscopic thermodynamic result of the optimal trade-off between the ability to encode information about the environment in the very features of the network and the ability to transmit information across it.
-While the former is quantified by the von Neumann network entropy and the latter by the Helmholtz free energy, the two are combined into a single quantity tantamount to the thermodynamic efficiency.
-We take as a case study an algorithm originally proposed for creating and searching trees of patterns under the assumption that, akin to living system, the algorithm seeks the best internal representation of the outer world, namely a real world dataset of patterns.
-Remarkably, the optimal thermodynamic efficiency overlaps with the optimal pattern-matching efficiency, defined in terms of accuracy and computational cost, whether the former is measured at the resolution scale where the network transits from exhibiting a specific heat with a single peak to a scale-invariant region.
-Moreover the critical hypothesis is investigated and empirical evidence suggests that the system benefits from the vicinity to the critical point whether it is faced with higher-complexity environments, while it does not when the environment is simpler.
-Keywords von Neumann entropy  $\cdot$ Specific heat  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning
+This work computationally investigates an algorithm that was originally proposed for creating and searching trees of patterns under the assumption that it seeks the most efficient internal representation of the external world, namely a real-world dataset of patterns.
+The algorithm produces a transition from star graph to regular 1d lattice, passing through irregular scale-invariant trees, under the effect of a single control parameter.
+The relatively novel information-theoretic tools of the von Neumann entropy and its derivative, the specific heat, together with the recent definition of thermodynamic quantities, such as work and heat, for networks are reviewed and applied to inspect the tree structures produced by the formation protocol.
+We observe some consistency between the optimal thermodynamic efficiency for networks and the optimal pattern-matching efficiency, defined in terms of the trade-off between accuracy and computational cost when the former is measured at the meso-resolution scale involved in the phase transition from star graph to scale-invariant network.
+We adduce the hypothesis of a richer frequency content at the meso-scale of the self-organized network to explain this observation.
+Moreover, empirical evidence suggests that the system benefits from the vicinity to the mentioned phase transition if it is faced with higher-complexity environments, while it does not when the environment is simpler.
+A phase portrait of the tree-generating algorithm is provided and multiple phase transitions are identified.
+Keywords Trees  $\cdot$ Networks  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning  $\cdot$ von Neumann entropy  $\cdot$ Specific heat
 1 Introduction
-Why complex structures emerge in a real world environment is a fascinating question which has not found any fully satisfactory answer at the point of today.
-We aim at some computational empirical evidence suggesting that the emergence of complex structures is the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [1, 2].
-The latter is defined in terms of microscopic quantities thanks to the fairly newly developments in statistical physics of fields in network theory [2, 3, 4, 5].
-In living systems, natural selection is the way the information about the environment flows into the system’s genome which will than encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system.
-Thus the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6].
+Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer.
+In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2].
+The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics of fields in network theory [1, 3, 4, 5].
+In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system.
+Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6].
 How these traits are optimized to maximize the fitness to the environment is an open question.
@@
 Here we challenge this hypothesis.
-In our work a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
-We take as case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbour (1-NN) pattern recognition tasks and subsequently extended to k-NN [13].
-The algorithm employs a tree formation protocol where raw data are reorganized into tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing tree structure.
-At each step the new pattern is linked to either the currently parent pattern or to one of the k-NN children of the parent pattern.
-The choice is based on the relative distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern.
+In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
+We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13].
+The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network.
+At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern.
+The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern.
 The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.
@@
 For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out.
-For some intermediate value of $r$ a ragged tree structure is produced, while for $r\rightarrow 2$ a 1d-lattice is created.
-Remarkably, the algorithm produces exclusively trees, i.e.
-number of spanning trees is always 1, which are sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively.
-Using the metaphor introduced by Jacob for natural selection111 ”Natural selection does not work as an engineer works.
-It works like a tinkerer - who does not know exactly what he is going to produce but uses whatever he finds around him whether it be pieces of string, fragments of wood, or old cardboards; in short works like a tinkerer who uses everything at his disposal to produce some kind of workable object.”  [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a distance based rule.
-Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
-The Portegys algorithm has the desirable properties of being simple, explainable and subject to a single control parameter $r$.
-Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network.
-Most notably thermodynamic efficiency and pattern-matching efficiency demonstrate noticeable overlap.
-Figure 1: The three prototypical structures generated by the Portegys algorithm.
+For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created.
+The algorithm produces exclusively trees, i.e.
+loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ is the number of links and nodes respectively.
+Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
+Figure 1: Tree structures generated by the Portegys algorithm.
 The algorithm creates star-like networks up to 1d lattice when $r\rightarrow 2$.
@@
 Depicted networks are generated from a subset of 50 samples randomly extracted from the MNIST dataset with $r=0.05,0.95,1.50$ respectively.
-Classes are indicated by the 10 different colors in the color-bar.
+Classes are indicated by the 10 different colors in the color bar.
+The use of a hierarchical topology, onto which a high-dimensional dataset is mapped, offers insights into the structured disorder of the data [15] on the one hand, while on a practical perspective, it limits the number of comparisons needed in matching an unknown pattern to a prototypical example.
+Namely, the algorithm introduced by Portegys is equipped with a pattern-searching method described in Section 3.1.
+Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
+The Portegys algorithm has the desirable properties of being simple, explainable, and subject to a single control parameter $r$.
+Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning, and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network.
+Notably, we observe some consistency between the thermodynamic efficiency and pattern-matching efficiency that we try to explain in terms of average frequency content in the self-organized network structure.
 2 Theoretical framework
@@
 Information spreading over networked structures is governed by the Laplacian matrix [16] and its spectrum of eigenvalues encodes many relevant topological properties of the graph [17, 18].
-Let us consider an undirected and unweighted network $G(V,E)$ of order $|V|=N$ with number of links equal to $|E|=m$.
+Let us consider an undirected and unweighted network $G(V,E)$ of order $|V|=N$ with the number of links equal to $|E|=m$.
 The network Laplacian is defined as $\hat{L}=\hat{D}-\hat{A}$, where $\hat{D}$ and $\hat{A}$ are the diagonal matrix of the node degrees and the adjacency matrix respectively.
 The Laplacian operator can be used to define a density operator [5], where the Laplacian takes the role of the Hamiltonian of statistical physics [19] as in Equation 1.
-Given an initial state of the network encoding the concentration of quantity, e.g.
+Given an initial state of the network encoding the concentration of a quantity, e.g.
 information, in one or more nodes, $\underline{s}(0)$, the state at time $\tau$ is given by $\underline{s}(\tau)=e^{-\tau\hat{L}}\underline{s}(0)$.
@@
 Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$.
-This relation remarks that the eigenvectors of the Laplacian are akin to the wave-vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the ”wave length” $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3].
-Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed defining a probability measure over the accessible modes decomposed by the spectral theorem.
+This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the "wavelength" $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3].
+Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed by defining a probability measure over the accessible modes decomposed by the spectral theorem.
 Such canonical ensemble [5, 4, 22] is encoded into the density matrix:
-$\hat{\rho}_{\tau}=\frac{e^{-\tau\hat{L}}}{Z}$ (1)
+$\hat{\rho}(\tau)=\frac{e^{-\tau\hat{L}}}{Z}$ (1)
 The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$.
@@
 Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
-Finally, the network von Neumann entropy can be introduced [5] as a multi-scale descriptor of the network structure:
+Finally, the von Neumann spectral entropy can be introduced [5] as a multi-scale descriptor of the network structure:
 $S(\hat{\rho}(\tau))=-\sum_{i=1}^{N}\mu_{i}(\tau)\log\mu_{i}(\tau)$ (2)
@@
 The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes.
-Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [21] as illustrated in Figure 2.
-Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the homogeneous phase where the information has propagated all over the network.
-The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/\log\tau$ it is possible to identify the characteristic resolution scales of the network [21, 3].
-For instance, a constant specific heat reflects the scale-invariance of the network where information is lost at a constant rate.
-Such a situation is typical of d-dimensional lattices and scale invariant networks with power law distributed spectrum.
-On the contrary, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network.
-The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network, see purple curve in Figure 1.
-Before the peak the informationally connected components are the single nodes and after the peak the informationally connected components components are the full-network revealing a single trivial-scale.
-From this considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 21].
-In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction.
+Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22].
+Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
+The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/d\log\tau$ it is possible to identify the characteristic resolution scales of the network [22, 3].
+For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network.
+The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network.
+Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale.
+On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate.
+Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3].
+The latter sentence makes for a definition for the scale-invariant networks [23].
+The specific heat plateau height, $c_{0}$, is also related to the spectral dimension of the network $d_{s}=2c_{0}$ [3, 23].
+From these considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 22].
+In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [19].
 To conclude, we can intuitively use the diffusion parameter $\tau$ to resolve the network topological structures at different resolution scales.
-Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the ”wavelengths” of its modes.
-Figure 2: Von Neumann entropy as a function of the diffusion time.
-Panel (a) depicts diffusion over two exemplary networks with analogous topological features to those depicted in panel (b): in blue a stochastic block model network exhibiting tripartite modular structure, and in purple a Erdös-Rényi random network with the same number of nodes of the blue one.
-For small $\tau\sim 0$, both networks exhibit the maximum spectral entropy value $\log N$, where $N$ is the number of nodes, reflecting the segregated phase.
-For $\tau\rightarrow\infty$, the spectral entropy converges to $\log C$, where $C$ is the number of connected components, thus reflecting the integrated phase.
-For intermediate values of $\tau$, the Erdös-Rényi network shows a single critical scale at which the network transitions from the segregated to the integrated phase.
-The position on the horizontal axis of the critical scale is related to the edge density of the network as well as the heterogeneity in the degree distribution.
-On the contrary, the blue network exhibits two critical resolution scales in the entropic transition.
-At smaller $\tau$, information initially diffuses within the modules ending into a plateau.
-A later slower diffusion process between the three modules completes the transition to the integrated phase.
-The insets show the density matrix $\hat{\rho}_{\tau}$ across the entropic transition on the blue curve.
-The tripartite modular structure of the network is shown by the emergence of the three blocks in the density matrix.
-Graphs depicted in (b) have only illustration purpose.
-Illustration from [22, 21, 23].
+Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the "wavelengths" of its modes.
 2.2 Thermodynamics of networks
-The statistical physics theory for networks introduced above naturally allows to view network phenomena from a new perspective and to bridge the gap to network macroscopic observables.
-To this regard network thermodynamic quantities as work, $W$, and heat, $Q$, can be defined in (at least mathematical) analogy to their thermodynamic counterparts [2].
+The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables.
+To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1].
 Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network.
-In turns, lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits.
+Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits.
 The entropic cost is given by:
-$Q=-\frac{dS}{\tau}=-\frac{\log N-S}{\tau}\leq 0$ (3)
-which resembles the heat dissipation in a thermodynamic process.
-On the other hand, it is obvious that adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [2, 4].
+$Q=-\frac{\delta S}{\tau}=-\frac{\log N-S}{\tau}\leq 0$ (3)
+which mathematically resembles the heat dissipation in a thermodynamic process.
+On the other hand, adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4].
 Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information.
@@
 Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
-$W=dF=F-F_{iso}=F+\frac{\log N}{\tau}\geq 0$ (4)
+$W=\delta F=F-F_{iso}=F+\frac{\log N}{\tau}\geq 0$ (4)
 which is tantamount to the work in the canonical ensemble.
@@
 A formation process can be depicted as the trade-off between these two necessities of the networked complex system: the need to maximize the information flow, while at the same time retaining an amount of structural entropy such that the network can encode and express sufficient diversity through its traits.
-Thus, to complete the analogy to thermodynamics, the trade-off between diversity of information and information flow is captured by the efficiency [2]:
+Thus, to complete the analogy to thermodynamics, the trade-off between the diversity of information and information flow is captured by the efficiency [1]:
 $\eta=\frac{W+Q}{W}=1-\frac{|Q|}{W}=\frac{U}{W}$ (5)
 where $0\leq\eta\leq 1$ and $U\equiv-\partial_{\tau}\log Z\equiv\operatorname{Tr}[\hat{L}\hat{\rho}]$.
-The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness and heterogeneity [2].
-In our work we make use of this variational principle to explain the structural properties of a network that is required to encode the diversity expressed in the environment, where the environment is defined by a real-world dataset of patterns.
-We also make use of the specific heat to identify which is the resolution scale at which such variational principle is effective for heterogeneous networks constrained to a sparsity regime.
+The variational principle here exposed has been devised to macroscopically explain the sparsity and its empirical scaling law, as well as the emergence of topological features such as modularity, small-worldness, and, with partially less clear results, heterogeneity [1].
+In our work, we use this variational principle to explain the structural properties of the network subclass composed of heterogeneous hierarchical trees.
+An additional element is that these trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns.
+The specific heat is used here to identify which is the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime.
+For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=c_{0}/\tau$ [3, 23], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{c_{0}}{W}$.
 3 Methods
-Our results are evaluated over subsets of three different datasets: the NIST [25], the MNIST [26] and the FashionMNIST [27] datasets.
+Our results are evaluated over subsets of four different datasets: the NIST [25], the MNIST [26], the FashionMNIST [27], and the CIFAR10 [28] datasets.
 These datasets contain randomly collected instances of image samples out of 10 classes.
 It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure.
+Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work.
+In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32].
+In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each.
+The left axis is referred to the quantities in green and orange, while the right axis is referred to the multilayer-perceptron accuracy.
+We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
 The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms.
-The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and white images of the handwritten digits [0-9].
+The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and-white images of the handwritten digits [0-9].
 The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories.
-In all our experiments a subset of 2000 samples is extracted from the train-set and the test-set independently.
+The CIFAR-10 dataset contains $32\times 32$ color images belonging to 10 classes.
+In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently.
+Samples from the FashionMNIST and MNIST are always binarized, while the samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal.
 In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST.
-The 10 classes are equally represented both in the test and in training set.
-We limit to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
-The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] and where distance between patterns is the Euclidean distance.
+The 10 classes are equally represented both in the test and in the training set.
+The datasets used in our experiments exhibit different levels of complexity, see Figure 2.
+We estimate with multiple methods the dataset complexities on the same subset of 2000 training samples used in the rest of the manuscript.
+The first heuristic approach is based on the test accuracy of a small multilayer perceptron (2 layers with 128 units each).
+In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes.
+A simple but effective parameter to gauge this facet of complexity is the accuracy that a classifier (such as the MLP) can achieve in determining the true class label for a data point.
+Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset.
+The second method we use is the computation of the Lempel-Ziv complexity [29] of the aggregated training set.
+In this case, samples need to be binarized.
+The third method concerns the measure of the intrinsic dimension [15, 32] of the dataset using the Maximum Likelihood Estimation approach introduced in [30] with the correction proposed in [31].
+A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the "world" of our problem, for instance, handwritten digits, it spans a lower dimensional manifold.
+Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [15, 32].
+We set the hyper-parameter $n=3$.
+While in Section 4 we discuss the full range of $r$, in Section 5 we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
+The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] where the distance between patterns is the Euclidean distance.
+Figure 3: Search method based on relative distance.
+In (a) the growth process is illustrated.
+At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$.
+The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node.
+Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node $P$ and the new pattern.
+In (b) the search method based on relative distance is illustrated.
+Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$.
+The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$.
+In the depicted case, patterns closer to the search pattern $S$ are expected to be in the subtree of $C_{2}$.
+Adapted from [33].
 3.1 Pattern-matching efficiency
-After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to a test pattern is given using a first-best approach.
-The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list, while providing a considerable speed up due to the tree-based structure.
-The search process is an iterative process which begins with the computation of the distance between the test pattern and the root-node pattern scaled by $r$.
-The distance is then compared to the distance between the test pattern and all of the root node’s children.
-The first node with a pattern that has a smaller distance is than selected as the new root node, and the process starts over with the newly selected node as root.
-The search phase terminates when no node whose pattern has smaller distance to the test pattern than the current one can be found.
-It is natural to define an pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
+After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach.
+The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure.
+The growth process, described in Section 1 and depicted in panel (a) of Figure 3, ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels.
+This provides the possibility of using a search technique based on relative distances.
+The search method is depicted in panel (b) of Figure 3 and is an iterative process that begins at the root-node pattern.
+Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtree, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Figure 3.
+The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child $C_{i}$, collectively indicated by the gray circle $C^{\prime}_{i}$ in panel (b).
+Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties.
+If $d(S,C_{i})_{search}$ is less than zero then it is set to zero.
+In panel (b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
+It is natural to define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
 $\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
-the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$ where $N$ is the number of nodes.
-For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases as does the accuracy but at a different pace, thus, it is possible to identify an optimal trade-off between the gain in computational burden and the loss in performance, as depicted in panel (c) of Figure 3.
-Figure 3: Pattern-matching efficiency, (c), is defined as the trade-off between the classification accuracy, (a), and the logarithm of the number of distances computed to find the best match, (b).
-The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency.
+the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes.
+For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases, potentially decreasing the class-match accuracy.
+When defining the pattern-matching efficiency we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or terms of general resources.
+Hence, the efficiency $\theta$ quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in panel (c, f) of Figure 4 for the MNIST and CIFAR10 respectively.
+Note in panel (d) that the CIFAR10 dataset has a peak for accuracy which we interpret by saying that for complex datasets (see also the case of the FashionMNIST in Section S7) some network structures do possess the ability to better encode information of the "world" in their traits.
+We postpone the discussion of this finding to Sections 5 and 6.
+Figure 4: Pattern-matching efficiency in (c, f) measured over the NIST and CIFAR10 respectively.
+The efficiency $\theta$ is defined as the trade-off between the classification accuracy, (a, d), and the logarithm of the number of distances computed to find the best match, (b, e).
+The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency $\theta$.
 The shaded grey area designates the standard deviation.
-The purple shaded area in panel (c) indicates a tolerance $r$-domain of optimal efficiency.
-The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by yellow shaded area.
-Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST dataset.
-The 10 digit-classes are equally represented both in the train- and test-set.
+The purple shaded area in panel (c, f) indicates a tolerance $r$-domain of optimal efficiency.
+The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area.
+Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST and CIFAR10 datasets.
+The 10-digit classes are equally represented both in the train- and test sets.
+4 Phase portrait
+Figure 5: Multiple phase transitions in hierarchical tree networks.
+In (a) the specific heat $C$ across the range $r\in[0,2]$.
+At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat.
+The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line.
+For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot c_{0}$, where $c_{0}$ is the plateau height in (c).
+Note that at the transition point $d_{s}=4/3$.
+In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.25,~{}0.35,~{}0.40,~{}0.70,~{}0.90,~{}1.05,2.00\}$.
+Dashed blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively.
+In (d), (e) the modularity and the normalized diameter, $\hat{d}=d/m$, together with their respective variances (dashed lines) identify the transition points.
+The modularity resolves the transition involving the mesoscale structure, while the diameter resolves the transition at the larger network scale.
+The legend indicates the two network sizes $N=500,~{}1000$.
+In (f) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-scales, i.e.
+the number of nodes included in the scale-invariant region.
+Note the discontinuity at the transition from star to amorphous.
+The dotted line shows the peak of the variance at the transition point.
+In (g) the entropy surface is depicted and the entropy at the first and second specific heat peaks are illustrated, see legend.
+In (h) the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate.
+In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right are depicted illustrating the two phase transitions at the two resolution scales.
+All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
+The Portegys algorithm produces tree structures as those depicted in panel (b) of Figure 5 depending on the value of the control parameter $r$.
+Two structural phase transitions involving two different resolution scales can be identified (see panel (a)).
+The transition at the largest resolution scale can be described using as an order parameter a quantity that describes the large-scale network properties: $\hat{d}=d/m$, where $d$ is the diameter of the tree and $m$ is the number of edges equal to $m=N-1$.
+For hierarchical trees, another choice would be the normalized number of hierarchical levels $\hat{h}=h/(N-1)$ where $h$ is the number of hierarchical levels ranging from $h\in[0,N-1]$ (see illustration in panel (b)).
+Both quantities describe the "volume" expansion controlled by the parameter $r$.
+We chose the normalized diameter $\hat{d}$ and we identify this phase transition at $r\sim 1.05$ as indicated by the peak of the variance (dotted lines) in panel (e) for the two network sizes $N=500,~{}1000$ as indicated in the legend.
+The network structure transits from the amorphous regime to the regular 1d-lattice.
+Just below the transition point, a spectral dimension of $d_{s}=4/3=2c_{0}$ characteristic of random trees [34] is measured as indicated in (a) and shown in panels (c).
+Above the transition $d_{s}=1$, as it is expected for a 1d-lattice.
+Where $c_{0}$ is the specific heat plateau height.
+The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$.
+Here, the emergence of scale-invariant structural properties is identified by the appearance of a second peak in the specific heat and a flat plateau in between the two $C$-peaks, as shown in panels (a) and (c) of Figure 5.
+Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph.
+The emergence of a plateau $C=c_{0}>0$ for $r>0.35$ reflects the emergence of progressively larger scale-invariant structures that for $r\rightarrow 2$ would include the full network in the thermodynamic limit $N\rightarrow\infty$.
+This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies occur either in the vicinity of this transition border or are included in the scale-invariant regime.
+This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network.
+In fact, we observe in panel (e) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition.
+Nevertheless, a phase transition indeed exists and an appropriate order parameter needs to be employed to identify it.
+This phase change occurs at smaller resolution scales, specifically, the mesoscales involved in the structural transition from the star graph to the scale-invariant regime (see the bifurcation in the specific heat and the related quantities in panels (a,c,g,h)).
+Thus, we can identify this transition using as an order parameter the modularity [35] of the tree which we show in panel (e).
+We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36].
+We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor.
+This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
+In panel (f), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks counts the normalized (log)number of informationally connected nodes across the $\tau$-scales where the entropy is consumed at a constant rate, i.e.
+across the specific heat plateau.
+In simple terms, we count the fraction of nodes included in the scale-invariant region.
+We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
+As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see panel (h)).
+For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase.
+Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].
 5 Results
-4.1 Thermodynamic and pattern-matching efficiency at the emergence of new C-peaks
-Figure 4 shows the coherence between algorithmic defined in Section 3 and thermodynamic efficiency in Equation 5 for the three investigated datasets.
-In panels (a)-(c) the pattern-matching efficiency, $\theta$, is depicted.
-The vertical red-dashed line indicates the $r$ value for the optimal algorithmic trade-off, while the shaded purple region indicates a $r$-domain tolerance of the optimal pattern-matching efficiency.
+Figure 6 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
+In panels (a)-(d) the pattern-matching efficiency, $\theta$, is depicted.
+The vertical red-dashed line indicates the $r$ value for the optimal trade-off, while the shaded purple region indicates a $r$-domain tolerance.
 The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
-In panels (d)-(f) the thermodynamic efficiency, $\eta$, shows remarkable overlap with the pattern-matching efficiency $\theta$ in terms of optimal $r$.
-See the $r$-range spanned by the purple-shaded areas in panels (a)-(c) and the vertical dashed lines in panels (d)-(f).
-The overlap between the two efficiencies suggests that, whether thermodynamic or algorithmic, the two objects are the same thing just viewed from different angles.
-See the next Section for the discussion on the double peak of the $\eta$ curves for the MNIST.
+In panels (e)-(h) the thermodynamic efficiency, $\eta$, is depicted and, for the case of the NIST and the FashionMNIST a notable consistency in the global maxima is observed with the pattern-matching efficiency $\theta$ in terms of optimal $r$.
+See the $r$-range spanned by the purple-shaded areas in panels (a)-(d) and the vertical dashed lines in panels (e)-(h).
+This consistency in the maxima of the two efficiencies suggests that thermodynamic or pattern-matching might be related to the context of this algorithm.
+In the Section 6 we discuss this matter.
 Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph.
-Thus, the emergence of structure is reflected by the arise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (e)-(i).
-Such transition is also depicted in Figure 5 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and for $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (g)-(i) of Figure 4.
-Before the bifurcation the two curves overlap and the single peak in the specific heat identifies a single trivial resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph.
-The emergence of a plateau $C=c_{0}>0$ for $r\rightarrow 2$ reflects the emergence of a progressively larger scale invariant structure.
-For $r\rightarrow 2$ we expect the plateau height to converge to the theoretical value $C\rightarrow c_{0}=1/2$ of the 1d-lattice passing by $c_{0}=4/6$ for random trees notoriously related to the spectral dimension $d=2c_{0}$ [3].
-See also Section S4 illustrating the spectral entropy and specific heat curves for the three datasets independently depicted for each $r$-value.
-Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the structural transition.
-Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5.
-In the seminal work [2], it is indicated to compute the thermodynamics efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
-the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium according to the authors, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
-Nevertheless, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$.
-Panels (a)-(c) of Figure 5 show the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves discussed above.
-The key difference is in the use of the ensemble average which employs all the eigenvalues to compute the specific heat.
-Thus, it is crucial to identify which is the $\tau$-scale which resolves the phase transition from the single trivial scale of the star graph to more complex topological traits of the amorphous network depicted in Figure 1.
-Such resolution scale is the relevant scale for the computation of the efficiency $\eta$ and defines an ”ultra-violet” cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
-In Section S3 the efficiency, $\eta$, is computed in the range interval indicated by the diffusion time as defined in [2] $\tau_{diff}=1/\lambda_{2}$, i.e.
+Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=c_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset.
+Such transition is also depicted in Figure 7 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 6.
+See also Section S4 illustrating the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
+Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Section 4.
+Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (i)-(l) of Figure 6 and in panels (a)-(c) of Figure 7.
+In the seminal work [1], it is indicated to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
+the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
+Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$.
+As shown in panel (h) of Figure 5 and in panels (a)-(d) of Figure 7 where the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves are illustrated.
+The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat.
+Thus, it is crucial to identify which is the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in panel (b) of Figure 5.
+Such resolution scale is the scale we consider of significant importance for the computation of the efficiency $\eta$ across the transition and defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
+In Section S3 the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e.
 the red curves depicted in panels (a).
-Whether computed at those scales, $\eta$ shows a change in concavity and no consistent relation with the algorithmic accuracy.
-Moreover note in Figure 5 the slow down of diffusion indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition.
-We remind that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
+Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
+Moreover note in Figure 7 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition.
+We note that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
 Figure 6: Thermodynamic & pattern-matching efficiency.
-In (a)-(c) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the three investigated datasets of increasing complexity (see arrow at the bottom of the figure).
+In (a)-(d) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure).
 The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line.
-In (d)-(f) the optimal thermodynamic efficiency, $\eta$, as a function of $r$ overlaps with the pattern-matching efficiency.
-Vertical dashed-lines indicate the global maxima of the efficiencies.
-Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (g)-(i).
+In (e)-(h) the optimal thermodynamic efficiency, $\eta$, as a function of $r$.
+In panels (f) and (h) a noticeable cohesion of the global maxima of the two efficiencies is observed.
+Vertical dashed lines indicate the global maxima of the efficiencies.
+Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l).
 The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure.
-Note that for datasets of higher complexity the highest efficiency, both thermodynamic and algorithmic, are posited at the critical value of the structural phase transition illustrated by the specific heat in panels (g)-(i).
-This is in line with the argument that states that a fairly simple dataset does not benefit from the critical point.
+Note that for datasets of higher complexity the highest efficiency, both thermodynamic and pattern-matching, tends to be posited closer to the structural phase transition to scale-invariant illustrated by the specific heat in panels (i)-(l).
+This is in line with the argument that states that a fairly simple dataset does not benefit from the vicinity of the phase transition.
 Quantities are averaged over 50 independent simulations.
 Figure 7: Role of the ensemble average on the estimation of the $\tau$-scale.
-Panels (a)-(c) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$.
+Panels (a)-(d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$.
 $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 6.
-$\tau_{diff}$ is the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$.
-Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (g)-(i) of Figure 4.
+$\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$.
+Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i)-(l) of Figure 6.
 At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$.
@@
 The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
-4.2 Performance at criticality
-As previously discussed, the $\tau$-scale where the bifurcation in panels (e)-(i) of Figure 4 and in panels (a)-(c) of Figure 5 shows up indicates the scale involved in the structural transition and it is the resolution scale where the thermodynamic efficiency overlaps with the pattern-matching efficiency as a function of $r$.
+5.1 Performance at the emergence of scale-invariance
+As previously discussed in Section 4, in panels (i)-(l) of Figure 6 and in panels (a)-(d) of Figure 7, the $\tau$-scale where the bifurcation shows up indicates the scale involved in the structural transition, and it is the resolution scale where the thermodynamic efficiency for some of the datasets displays consistency with the pattern-matching efficiency as a function of $r$.
 Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
-To discuss this empirical result we turn our attention to the panels (a)-(c) in Figure 4 and panels (a)-(c) in Figure 5.
-While for the FashionMNIST dataset there is a rough correspondence between the $r\sim 0.30$ where the structural transition emerges and the $r$-range $r\in(0.30,0.50)$ where the global maximum of the efficiencies is observed, for the MNIST and NIST datasets the thermodynamic and algorithmic efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.35$ and $r\sim 0.45$.
-The term ”far away” is quantified by the purple-shaded areas Figure 4.
-We rather observe in panels (b), (c), (e) and (f) of Figure 4 that the maximum efficiencies are posited where the $\tau^{(2)}_{peak}$ have a discontinuous jump in panels (b), (c) of Figure 5 at $r\sim 0.60$ for the MNIST and at $r\sim 0.70$ for the NIST.
-The argument holds also for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
-The jump for the MNIST curve in Figure 5 is such that the two $\tau^{(1,2)}_{peak}$ curves overlap.
-We mention that this effect is mitigated for larger datasets, i.e.
+To discuss this empirical result we turn our attention to the panels (a)-(d) and (i)-(l) in Figure 6 and panels (a)-(d) in Figure 7.
+For the CIFAR10 and FashionMNIST datasets there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the global maximum of the efficiencies is observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively for the two datasets.
+On the other hand, for the MNIST and NIST datasets the thermodynamic and pattern-matching efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$.
+The term "far away" is quantified by the purple-shaded areas in panels (a)-(d) of Figure 6.
+The argument also holds for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
+For the MNIST case, we observe that a third large peak emerges in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 6 and the entropy curves depicted in Figure S6).
+This effect is only mitigated for larger datasets, i.e.
 larger networks, as we show in Section S2 of the Supplementary.
-The jump is caused by the presence of a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Figure 4 and the entropy curves depicted in Figure S6).
-Moreover, for the MNIST case, the jump is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$.
-Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 4(b), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 4(e).
-Nevertheless, we stress once again that we did not compute the thermodynamic efficiency at the corresponding $\tau$ scale where discontinuity of the $\tau^{(2)}_{peak}$ curve appears, but we rather measured $\eta$ at the $\tau$ scale indicated by the bifurcation.
-To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or father away, we turn to the argument, already present in the literature [28, 10, 29], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only weather the task requires a high level of complexity.
-While for the FashionMNIST, a notoriously more complex dataset when compared to both the MNIST and NIST, the optimal value for the efficiency is posited in the proximity of the critical value of $r\sim 0.30$, the optimal values of $r$ for the NIST and MNIST datasets are evidently far away from their respective transition values.
-The MNIST, which is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$.
-While the global maximum in Figure 4(e) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 5, the local maxima is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation.
-This observation is in line with the hypothesis that the system takes advantage from the critical behaviour only when a fairly complex task is presented to it and in line with other observations discussed in the literature [28, 10].
-Thus, it stresses the importance of dataset under analyses.
+This intermediate peak is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$.
+Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 6(c), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 6(g).
+We stress that also in this case, we measured $\eta$ at the $\tau$ scale indicated by the bifurcation $\tau\sim 10$.
+To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or farther away, we turn to the argument, already present in the literature [37, 10, 38], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only if a high level of complexity is required by the task, see Figure 2.
+While for the CIFAR10 and FashionMNIST, more complex datasets when compared to both the MNIST and NIST, their respective optimal value for the efficiency is posited in the proximity of the phase transition, the optimal values of $r$ for the NIST and MNIST datasets are far away from their respective transition values.
+The MNIST, that is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$.
+While the global maximum in Figure 6(g) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 7, the local maximum is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation.
+This observation is in line with the hypothesis that the system takes advantage of the critical behavior of a certain resolution region only when a fairly complex task is presented to it and in line with other observations discussed in the literature [37, 10].
+Thus, it stresses the importance of the dataset under analysis and its disordered distribution of points in a high-dimensional feature space.
 […]
@@
 class-matching.
-By using the recently developed theoretical tools of statistical physics of fields for networks [5, 4, 34, 21, 3, 2], we have quantified the diversity in the traits of the network with the von Neumann entropy and the transmission flow of information with the Helmholtz free energy.
-As fastness of information flow and sufficiently rich representation of the outer world are two competing capabilities, the formation process of a complex networked structure is interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [2].
-We have shown that such a thermodynamic efficiency strikingly overlaps with the pattern-matching efficiency defined in terms of accuracy and computational costs, suggesting that the two can be regarded as the same thing but viewed from different angles.
-Notably, we have shown that the thermodynamic efficiency needs to be computed at a specific resolution scale $\tau$.
-We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial-structure, that is a single specific heat peak, to emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
-Such $\tau$-resolution scale is akin to an ”ultra-violet” cut-off delimiting the frequency of the network structure across the heterogeneity-transition from star graph to 1d-lattice trough amorphous.
-While the transition point reflects the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency.
-We explain this with the statement that a system benefits from the critical point only when it confronts with environments of higher complexity, while it does not benefit from it when the environmental complexity is lower.
-This is elucidated by the use of three different datasets of increasing complexity and supported by other works in the literature [28, 10, 29].
+The recently developed theoretical tools of statistical physics of fields for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy.
+As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
+In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure.
+The tree-formation undergoes two phase transitions at two different scales: a meso-scale with the emergence of a approximately scale-invariant structural behaviour that progressively involves a larger scales, and a large-scale transition from the irregular scale-invariant behaviour to the regular-scale invariant 1d lattice.
+Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different level of complexity.
+When the dataset was more difficult, an optimal efficiency to the pattern-searching task is found to be close to the transition to scale-invariant in the parameter space.
+Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range.
+Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an undirected process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
+In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$.
+For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$.
+It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4].
+They are not exactly alike as one indeed involves a directed process while the former involves an undirected process; still, they can be used as estimates of the "flow".
+We have discussed the denominator of the two efficiencies: $\eta=U/W$ and $\theta=\text{Accuracy}/\log N_{dist}$.
+We now discuss the numerator.
+For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
+The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping.
+Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one to one relation between the two.
+In the case of a network structure the parameters are the network traits themselves, which in the Fourier space become the “network frequencies” $\lambda_{i}$, i.e.
+the Laplacian eigenvalues.
+We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5.
+The internal energy $U$ is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale $\tau$, tantamount to the inverse temperature $\beta$ in statistical physics.
+Thus, the Gibbs average frequency content is quantified by $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$ the internal energy, where for scale-invariant networks $<\lambda>_{\tau}=c_{0}/\tau$ [3].
+The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given an external control parameter $r$.
+Notably, we have computed the thermodynamic efficiency at a specific resolution scale $\tau$.
+We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
+Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e.
+from star graph to amorphous.
+The operational $\tau$-scale employed roughly identifies modules in the network as at the $r$-bifurcation point we have observed a phase transition when employing the modularity as an order parameter.
+While the transition point to scale-invariant permits the identification of the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency.
+We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower.
+This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
+In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of solely the Portegys algorithm.
+Given a pre-trained tree network structure on a dataset that is unknown, it is possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the thermodynamic efficiency on the phase portrait without performing any test phase.
+As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, it posits itself close to the phase transition boundary at the cost of a slower pattern search.
+Interestingly, our analysis on hierarchical tree networks could be also applied, if further developed, to other topics in artificial intelligence.
+We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing.
+In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals.
+Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years.
+In the neuromorphic paradigm computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio.
+The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48].
+In materio tree-like structures are currently physically implemented.
+We mention tree structures produced at the critical point of a percolation transition [49, 50, 51] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
 Conflict of Interest Statement

### v4

@@
 In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2].
-The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics of fields in network theory [1, 3, 4, 5].
+The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics in network theory [1, 3, 4, 5].
 In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system.
@@
 The algorithm produces exclusively trees, i.e.
-loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ is the number of links and nodes respectively.
+loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively.
 Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
@@
 Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [20, 22].
-The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location at time $\tau$ [4].
+The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location after time $\tau$ [4].
 Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
 The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes.
-Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22].
+Thus the von Neumann entropy, $S$, as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22].
 Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
-The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/d\log\tau$ it is possible to identify the characteristic resolution scales of the network [22, 3].
+The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [22, 3].
 For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network.
@@
 To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1].
-Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network.
+Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges diminishes the network’s entropy.
 Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits.
@@
 which mathematically resembles the heat dissipation in a thermodynamic process.
-On the other hand, adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4].
+On the other hand, adding edges increases the gain in information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4].
 Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information.
@@
 In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each.
-The left axis is referred to the quantities in green and orange, while the right axis is referred to the multilayer-perceptron accuracy.
+The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy.
 We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
@@
 Figure 5: Multiple phase transitions in hierarchical tree networks.
-In (a) the specific heat $C$ across the range $r\in[0,2]$.
+In (a) the specific heat, $C$, across the range $r\in[0,2]$.
 At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat.
@@
 For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot C_{0}$, where $C_{0}$ is the plateau height in (c).
-Note that at the transition point $d_{s}=4/3$.
-In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.25,~{}0.35,~{}0.40,~{}0.70,~{}0.90,~{}1.05,2.00\}$.
-Dashed blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively.
-In (d), (e) the modularity and the normalized diameter, $\hat{d}=d/m$, together with their respective variances (dashed lines) identify the transition points.
-The modularity resolves the transition involving the mesoscale structure, while the diameter resolves the transition at the larger network scale.
-The legend indicates the two network sizes $N=500,~{}1000$.
-In (f) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-scales, i.e.
+Note that at $r\sim 0.8$ the spectral dimension is $d_{s}=4/3$ as expected for random trees with finite variance in the node degree.
+In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.35,~{}0.40,~{}0.70,~{}0.80,~{}0.90,2.00\}$.
+Vertical blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively.
+In (d), the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate.
+In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right.
+All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations.
+Figure 6: Order parameters.
+In (a), the modularity, $Q$, resolves the transition involving the mesoscale structure.
+In (b), the normalized diameter, $\hat{d}=d/m$ resolves the large-scale transition.
+Their respective variances (dashed lines) identify the transition points.
+In (c), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-domain, i.e.
 the number of nodes included in the scale-invariant region.
 Note the discontinuity at the transition from star to amorphous.
-The dotted line shows the peak of the variance at the transition point.
-In (g) the entropy surface is depicted and the entropy at the first and second specific heat peaks are illustrated, see legend.
-In (h) the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate.
-In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right are depicted illustrating the two phase transitions at the two resolution scales.
-All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
-The Portegys algorithm produces tree structures as those depicted in panel (b) of Figure 5 depending on the value of the control parameter $r$.
-Two structural phase transitions involving two different resolution scales can be identified (see panel (a)).
-The transition at the largest resolution scale can be described using as an order parameter a quantity that describes the large-scale network properties: $\hat{d}=d/m$, where $d$ is the diameter of the tree and $m$ is the number of edges equal to $m=N-1$.
-For hierarchical trees, another choice would be the normalized number of hierarchical levels $\hat{h}=h/(N-1)$ where $h$ is the number of hierarchical levels ranging from $h\in[0,N-1]$ (see illustration in panel (b)).
-Both quantities describe the "volume" expansion controlled by the parameter $r$.
-We chose the normalized diameter $\hat{d}$ and we identify this phase transition at $r\sim 1.05$ as indicated by the peak of the variance (dotted lines) in panel (e) for the two network sizes $N=500,~{}1000$ as indicated in the legend.
-The network structure transits from the amorphous regime to the regular 1d-lattice.
-Just below the transition point, a spectral dimension of $d_{s}=4/3=2c_{0}$ characteristic of random trees [34] is measured as indicated in (a) and shown in panels (c).
-Above the transition $d_{s}=1$, as it is expected for a 1d-lattice.
-Where $c_{0}$ is the specific heat plateau height.
+Curves in (a-b) are realized for networks of $N=500,~{}1000$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
+Curves in (c) are for networks of $N=500$.
+[…]
 The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$.
-Here, the emergence of scale-invariant structural properties is identified by the appearance of a second peak in the specific heat and a flat plateau in between the two $C$-peaks, as shown in panels (a) and (c) of Figure 5.
+Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in (a) and (c) of Figure 5.
+For $r\rightarrow 2$, the plateau would include the full 1d-lattice in the thermodynamic limit $N\rightarrow\infty$.
 Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph.
-The emergence of a plateau $C=c_{0}>0$ for $r>0.35$ reflects the emergence of progressively larger scale-invariant structures that for $r\rightarrow 2$ would include the full network in the thermodynamic limit $N\rightarrow\infty$.
-This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies occur either in the vicinity of this transition border or are included in the scale-invariant regime.
+After the phase transition, a second peak of the specific heat emerges at larger $\tau$ scales.
+This second peak of $C$ in Figure 5.(c) identifies the scale at which the information is integrated between the distinct modules.
+In other words, the second peak reflects the "degree" of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes.
+Note that we exclude both the clusters, i.e.
+triangles, and higher-order cycles from our discussion as the tree networks we address are purely loop-less.
+This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies always occur in the scale-invariant region of the phase plane, and in some cases, they occur in the vicinity of this transition border $r\sim 0.35$.
 This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network.
-In fact, we observe in panel (e) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition.
-Nevertheless, a phase transition indeed exists and an appropriate order parameter needs to be employed to identify it.
-This phase change occurs at smaller resolution scales, specifically, the mesoscales involved in the structural transition from the star graph to the scale-invariant regime (see the bifurcation in the specific heat and the related quantities in panels (a,c,g,h)).
-Thus, we can identify this transition using as an order parameter the modularity [35] of the tree which we show in panel (e).
+Indeed, we observe in Figure 6.(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition.
+Nevertheless, a phase transition exists and an appropriate order parameter needs to be employed to identify it.
+This phase change occurs at smaller resolution scales with respect to the first phase transition we discussed.
+Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation in the specific heat in Figure 5.(a),(c)).
+Thus, we can identify this transition using as an order parameter the modularity [35], $Q$, of the tree which we show in panel (a) of Figure 6.
 We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36].
@@
 This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
-In panel (f), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks counts the normalized (log)number of informationally connected nodes across the $\tau$-scales where the entropy is consumed at a constant rate, i.e.
-across the specific heat plateau.
+As a second possible order parameter, for the mesoscale phase-transition, we show in Figure 6.(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks.
+This order parameter counts the normalized (log)number of informationally connected nodes across the $\tau$-scales of the specific heat plateau.
 In simple terms, we count the fraction of nodes included in the scale-invariant region.
 We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
-As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see panel (h)).
+As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see Figure 5.(d)).
 For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase.
@@
 Figure 7 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
-In panels (a)-(d) the pattern-matching efficiency, $\theta$, is depicted.
-The vertical red-dashed line indicates the $r$ value for the optimal trade-off, while the shaded purple region indicates a $r$-domain tolerance.
-The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
-In panels (e)-(h) the thermodynamic efficiency, $\eta$, is depicted and, for the case of the NIST and the FashionMNIST a notable consistency in the global maxima is observed with the pattern-matching efficiency $\theta$ in terms of optimal $r$.
-See the $r$-range spanned by the purple-shaded areas in panels (a)-(d) and the vertical dashed lines in panels (e)-(h).
-This consistency in the maxima of the two efficiencies suggests that thermodynamic or pattern-matching might be related to the context of this algorithm.
+Panels (a)-(d) depict the thermodynamic efficiency, $\eta$, and panels (e)-(h) depict the pattern-matching efficiency, $\theta$.
+The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in (e)-(h) indicate a $r$-domain tolerance for $\theta$.
+The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
+Note the consistency in the global maxima of $\theta$ and $\eta$ in terms of optimal $r$ for the NIST and the FashionMNIST.
+See the $r$-range spanned by the purple-shaded areas in panels (e)-(h) and the vertical dashed lines in panels (a)-(d).
+This consistency in the maxima suggests that the two efficiencies might be related, for the specific context of this algorithm.
 In the Section 6 we discuss this matter.
-Similar to what would be the transition from vapor/liquid to an amorphous solid, at a critical value of $r$ topological complexity increases from the trivial traits of a star graph.
+The topological complexity of the networks increases at a critical $r$ from the trivial traits of a star graph to the fractal structure of the amorphous network.
 Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset.
 Such transition is also depicted in Figure 8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 7.
-See also Section S4 illustrating the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
-Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Section 4.
-Such critical resolution scale is identified by the bifurcation of the specific heat peaks depicted in panels (i)-(l) of Figure 6 and in panels (a)-(c) of Figure 7.
-In the seminal work [1], it is indicated to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
-the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, as this provides the time-scale where the diffusion process has reached equilibrium, and then logarithmically divide $\tau_{diff}$ to progressively define the large propagation scales, the middle scales and the small scales.
+See also Section S4 which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
+We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 7).
+This is an intrinsic characteristic of the dataset, as it is only mitigated for larger networks, as shown in Section S2 of the Supplementary.
+This intermediate peak is sufficiently strong to be reflected on a local maximum of $\eta$, in Figure 7.(c).
+Although this local maximum in $\eta$ is not detected by $\theta$ in Figure 7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area when we compare $\eta$ and $\theta$ peaks in terms of $r$.
+This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Figure 7.(c).
+Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Sections 4 and 6.
+Such critical resolution scale is identified by the bifurcation of the specific heat peaks in panels (i)-(l) of Figure 7 and in panels (a)-(d) of Figure 8.
+In the seminal work [1], it is suggested to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
+the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales.
+In fact, this eigenvalue should provide the time scale where the diffusion process has reached equilibrium.
 Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$.
-As shown in panel (h) of Figure 5 and in panels (a)-(d) of Figure 7 where the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves are illustrated.
+Panels (a)-(d) of Figure 8 depict the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves.
 The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat.
-Thus, it is crucial to identify which is the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in panel (b) of Figure 5.
-Such resolution scale is the scale we consider of significant importance for the computation of the efficiency $\eta$ across the transition and defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
-In Section S3 the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e.
+It is crucial to identify the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in Figure 5.(b).
+We consider such resolution scale of significant importance for the computation of the efficiency $\eta$ across the transition as it defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
+In Section S3 of the Supplementary Material the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e.
 the red curves depicted in panels (a).
 Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
-Moreover note in Figure 7 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$, we remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only which shapes the degree distribution as the number of nodes and edges is fixed across the transition.
-We note that the Portegy’s algorithm allows us to stay in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
+Finally, note in Figure 8 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$.
+We remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only, and the latter shapes the degree distribution only, as the number of nodes and edges is fixed across the transition.
+We note that the Portegy’s algorithm stays in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
 Figure 7: Thermodynamic & pattern-matching efficiency.
-In (a)-(d) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure).
+In (a)-(d) the thermodynamic efficiency, $\eta$, as a function of $r$.
+The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l).
+The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant "length" scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure.
+In (e)-(h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure).
 The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line.
-In (e)-(h) the optimal thermodynamic efficiency, $\eta$, as a function of $r$.
-In panels (f) and (h) a noticeable cohesion of the global maxima of the two efficiencies is observed.
+In panels (b)-(f) and (d)-(h) a noticeable cohesion of the peaks of the two efficiencies is observed.
 Vertical dashed lines indicate the global maxima of the efficiencies.
-Remarkably, the thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l).
-The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant length scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=c_{0}>0$, reflecting a roughly scale-invariant structure.
-Note that for datasets of higher complexity the highest efficiency, both thermodynamic and pattern-matching, tends to be posited closer to the structural phase transition to scale-invariant illustrated by the specific heat in panels (i)-(l).
-This is in line with the argument that states that a fairly simple dataset does not benefit from the vicinity of the phase transition.
+The efficiency peaks, both $\eta$ and $\theta$, are always in the amorphous region of the phase plane.
+For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star-graph to the amorphous region.
+This agrees with the argument that a fairly simple dataset does not benefit from the vicinity of the phase transition.
 Quantities are averaged over 50 independent simulations.
@@
 5.1 Performance at the emergence of scale-invariance
-As previously discussed in Section 4, in panels (i)-(l) of Figure 6 and in panels (a)-(d) of Figure 7, the $\tau$-scale where the bifurcation shows up indicates the scale involved in the structural transition, and it is the resolution scale where the thermodynamic efficiency for some of the datasets displays consistency with the pattern-matching efficiency as a function of $r$.
-Nevertheless, the critical value of $r$ where such structural transition appears is not necessarily the value where the efficiencies have the global maximum.
-To discuss this empirical result we turn our attention to the panels (a)-(d) and (i)-(l) in Figure 6 and panels (a)-(d) in Figure 7.
-For the CIFAR10 and FashionMNIST datasets there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the global maximum of the efficiencies is observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively for the two datasets.
-On the other hand, for the MNIST and NIST datasets the thermodynamic and pattern-matching efficiencies show a maximum in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$.
-The term "far away" is quantified by the purple-shaded areas in panels (a)-(d) of Figure 6.
-The argument also holds for the MNIST if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
-For the MNIST case, we observe that a third large peak emerges in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 6 and the entropy curves depicted in Figure S6).
-This effect is only mitigated for larger datasets, i.e.
-larger networks, as we show in Section S2 of the Supplementary.
-This intermediate peak is sufficiently strong to be reflected on a local maximum in the efficiency, $\eta$.
-Even though this local maximum in $\eta$ is not registered by the averaged pattern-matching efficiency $\theta$ in Figure 6(c), it is reflected in its standard deviation which allows us to include the $r$-range indicated by the shaded purple area, which corresponds to the $\eta$-peak-to-peak $r$-range in Figure 6(g).
-We stress that also in this case, we measured $\eta$ at the $\tau$ scale indicated by the bifurcation $\tau\sim 10$.
-To explain the diversity in the positioning of the maximum of the efficiencies, either in the vicinity of the transition point or farther away, we turn to the argument, already present in the literature [37, 10, 38], that indicates the necessity of tuning a system, or of a system to self-tune, at the border of a phase transition only if a high level of complexity is required by the task, see Figure 2.
-While for the CIFAR10 and FashionMNIST, more complex datasets when compared to both the MNIST and NIST, their respective optimal value for the efficiency is posited in the proximity of the phase transition, the optimal values of $r$ for the NIST and MNIST datasets are far away from their respective transition values.
-The MNIST, that is the intermediate dataset in terms of complexity, is an interesting case as it shows two maxima for the thermodynamic efficiency $\eta$.
-While the global maximum in Figure 6(g) is far away from the critical value of $r$ illustrated by the bifurcation in panel (b) of Figure 7, the local maximum is at $r\sim 0.35-0.40$ which corresponds to the critical value of $r$ indicated by the bifurcation.
-This observation is in line with the hypothesis that the system takes advantage of the critical behavior of a certain resolution region only when a fairly complex task is presented to it and in line with other observations discussed in the literature [37, 10].
-Thus, it stresses the importance of the dataset under analysis and its disordered distribution of points in a high-dimensional feature space.
-[…]
+In this section, we discuss the relation between the location of the pattern matching efficiency, $\theta$, on the phase plane of the algorithm and the complexity of the dataset, see Figure 2.
+We turn our attention to the panels (e)-(h) and (i)-(l) in Figure 7 and panels (a)-(d) in Figure 8.
+For the CIFAR10 and FashionMNIST datasets, there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the peaks of each of the two $\theta$ efficiencies are observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively.
+Note that for the FashionMNIST, also the thermodynamic $\eta$ peaks in the same range of $r$.
+On the other hand, for the MNIST and NIST datasets, $\theta$ peaks in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$.
+The term "far away" is quantified by the purple-shaded areas in panels (e)-(h) of Figure 7.
+Note the correspondence of the peaks of $\eta$ and $\theta$ for the two datasets in terms of $r$.
+The latter observation on the correspondence of $\theta$ and $\eta$ also holds for the MNIST, if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
+To explain the varying locations of peak efficiencies, whether near the transition border or farther from it, we refer to an argument in the literature [37, 10, 38].
+A system needs tuning, or self-tuning, at the edge of a phase transition only when the task demands a high level of complexity, specifically, this was investigated in the context of second-order phase transitions [37, 10].
+For the more complex datasets CIFAR10 and FashionMNIST, the optimal $\theta$ values are closer to the transition border.
+In contrast, the optimal values for the simpler NIST and MNIST datasets are located farther from their respective transition points.
+This observation also holds for the thermodynamic $\eta$ for the FashionMNIST, MNIST, and NIST.
+Moreover, the MNIST, an intermediate dataset in terms of complexity, is an interesting case.
+It shows two maxima for the thermodynamic efficiency, $\eta$.
+The global maximum in Figure 7(c) is far away from the critical value of $r$ illustrated by the bifurcation in panel (c) of Figure 8, while the local maximum is at $r\sim 0.35-0.40$, which corresponds to the critical value of $r$ indicated by the bifurcation.
+These observations align with the hypothesis that the system exploits its proximity to the phase transition boundary only when presented with a relatively complex task.
+The overall discussion highlights the importance of the dataset under analysis, particularly its disordered distribution of points within a high-dimensional feature space.
+Finally, we observe that while the value of $r$ where a structural transition appears is dependent on the dataset, i.e.
+$r=0.50$ for the NIST, $r=0.40$ for the MNIST, $r=0.35$ for the FashionMNIST and $r=0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is roughly the same for all the four datasets.
+Interestingly, the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary Section S2).
+We speculate about the possibility that this is due to the constraint on sparsity as it is an invariant feature over the cases investigated in this work.
 6 Discussion
@@
 In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset - they are posed in contact with.
-Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or in the very network-features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed-way as in artificial neural networks where the engineer designs a complicated distribution through the definition of an architecture and the training phase fills in the parameters in order to obtain the desired function-approximation machine to achieve statistical generalization [41].
-When confronted to real world environment, in order to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42].
-Moreover the same system needs to cope with cost of sending information across its structure which can cause an enhanced exploitation of resources.
-In the context of machine-learning, cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g.
+Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed way as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine to achieve statistical generalization [41].
+When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42].
+Moreover, the same system needs to cope with the cost of sending information across its structure which can cause an enhanced exploitation of resources.
+In the context of machine learning, the cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g.
 class-matching.
-The recently developed theoretical tools of statistical physics of fields for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy.
+The recently developed theoretical tools of statistical physics for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy.
 As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
 In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure.
-The tree-formation undergoes two phase transitions at two different scales: a meso-scale with the emergence of a approximately scale-invariant structural behaviour that progressively involves a larger scales, and a large-scale transition from the irregular scale-invariant behaviour to the regular-scale invariant 1d lattice.
-Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different level of complexity.
-When the dataset was more difficult, an optimal efficiency to the pattern-searching task is found to be close to the transition to scale-invariant in the parameter space.
+The tree formation undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
+Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different levels of complexity.
+The optimal pattern-matching efficiency was always found in the amorphous regime, Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star-graph and the scale-invariant regimes.
 Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range.
-Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an undirected process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
+Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an ergodic process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
 In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$.
@@
 It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4].
-They are not exactly alike as one indeed involves a directed process while the former involves an undirected process; still, they can be used as estimates of the "flow".
-We have discussed the denominator of the two efficiencies: $\eta=U/W$ and $\theta=\text{Accuracy}/\log N_{dist}$.
+They are not exactly alike as only the latter is an ergodic process; still, they can be used as estimates of the "flow".
+We have discussed the denominator of the two efficiencies: $\eta$ and $\theta$.
 We now discuss the numerator.
@@
 The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping.
-Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one to one relation between the two.
-In the case of a network structure the parameters are the network traits themselves, which in the Fourier space become the “network frequencies” $\lambda_{i}$, i.e.
+Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one-to-one relation between the two.
+In the case of a network structure, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e.
 the Laplacian eigenvalues.
 We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5.
-The internal energy $U$ is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale $\tau$, tantamount to the inverse temperature $\beta$ in statistical physics.
-Thus, the Gibbs average frequency content is quantified by $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$ the internal energy, where for scale-invariant networks $<\lambda>_{\tau}=c_{0}/\tau$ [3].
-The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given an external control parameter $r$.
-Notably, we have computed the thermodynamic efficiency at a specific resolution scale $\tau$.
+The internal energy, $U$, is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale, $\tau$, tantamount to the inverse temperature, $\beta$, in statistical physics.
+Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau$ [3].
+The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given the value of the external control parameter, $r$.
+Notably, we have computed the thermodynamic efficiency at a specific resolution scale, $\tau$.
 We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
@@
 from star graph to amorphous.
-The operational $\tau$-scale employed roughly identifies modules in the network as at the $r$-bifurcation point we have observed a phase transition when employing the modularity as an order parameter.
-While the transition point to scale-invariant permits the identification of the network resolution scale where the thermodynamic efficiency needs to be measured, the corresponding transition value of the control parameter $r$ does not coincide with the global maximum in efficiency.
+The operational $\tau$-scale employed roughly identifies the scale where modules emerge in the network.
+For this reason, we could gauge the phase transition employing the network modularity, $Q$, as an order parameter.
+While the transition to scale-invariant permits the identification of the network resolution scale, $\tau$, where the thermodynamic efficiency needs to be measured to observe a peak, the corresponding transition value of the control parameter, $r$, does not always coincide with the global maximum in the efficiency.
 We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower.
 This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
-In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of solely the Portegys algorithm.
-Given a pre-trained tree network structure on a dataset that is unknown, it is possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the thermodynamic efficiency on the phase portrait without performing any test phase.
-As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, it posits itself close to the phase transition boundary at the cost of a slower pattern search.
-Interestingly, our analysis on hierarchical tree networks could be also applied, if further developed, to other topics in artificial intelligence.
+In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of only the Portegys algorithm.
+Given a pre-trained tree network based upon an unknown dataset, it might be possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the peaks of the pattern-matching and thermodynamic efficiency on the phase plane without performing any test phase.
+As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, the peak of the pattern-matching efficiency might be at the phase transition boundary providing higher accuracy at the cost of a slower pattern search.
+Interestingly, our analysis of hierarchical tree networks could be applied, if further developed, to other topics in artificial intelligence.
 We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing.
@@
 Moreover, potential applications include the topic of neuromorphic computing [45, 46, 47] currently looking for a possible solution to the ever-increasing computational power demand of recent years.
-In the neuromorphic paradigm computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio.
+In the neuromorphic paradigm, computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio.
 The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [48].
-In materio tree-like structures are currently physically implemented.
-We mention tree structures produced at the critical point of a percolation transition [49, 50, 51] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
+Tree-like structures are currently physically implemented in materio.
+We mention tree structures produced at the critical point of a percolation transition [49, 50] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 51], but also fractal trees produced in the dielectric breakdown [52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
 Conflict of Interest Statement
@@
 Acknowledgments
-We are thankful to Jos van Goor for providing some of the scripts that enabled efficient grid-search across parameters and to Federico Balducci for the useful discussions.
+We are thankful to Jos van Goor for providing some of the scripts that enabled efficient grid search across parameters and to Federico Balducci for the useful discussions.
 Supplemental Data
-Supplementary Material include figures that are mentioned in the main text but are omitted from it to the purpose of clarity.
+The Supplementary Material includes figures referenced in the main text but omitted for clarity.
 Data Availability Statement

### v5

@@
 Abstract
-This work computationally investigates an algorithm that was originally proposed for creating and searching trees of patterns under the assumption that it seeks the most efficient internal representation of the external world, namely a real-world dataset of patterns.
-The algorithm produces a transition from star graph to regular 1d lattice, passing through irregular scale-invariant trees, under the effect of a single control parameter.
-The relatively novel information-theoretic tools of the von Neumann entropy and its derivative, the specific heat, together with the recent definition of thermodynamic quantities, such as work and heat, for networks are reviewed and applied to inspect the tree structures produced by the formation protocol.
-We observe some consistency between the optimal thermodynamic efficiency for networks and the optimal pattern-matching efficiency, defined in terms of the trade-off between accuracy and computational cost when the former is measured at the meso-resolution scale involved in the phase transition from star graph to scale-invariant network.
-We adduce the hypothesis of a richer frequency content at the meso-scale of the self-organized network to explain this observation.
-Moreover, empirical evidence suggests that the system benefits from the vicinity to the mentioned phase transition if it is faced with higher-complexity environments, while it does not when the environment is simpler.
-A phase portrait of the tree-generating algorithm is provided and multiple phase transitions are identified.
-Keywords Trees  $\cdot$ Networks  $\cdot$ Thermodynamics  $\cdot$ Statistical physics  $\cdot$ Machine Learning  $\cdot$ von Neumann entropy  $\cdot$ Specific heat
-1 Introduction
-Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer.
-In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2].
-The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics in network theory [1, 3, 4, 5].
-In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system.
-Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6].
+Hierarchical tree structures are common in many real-world systems, from tree roots and branches to neuronal dendrites and biologically inspired artificial neural networks, as well as in technological networks for organizing and searching complex datasets of high-dimensional patterns.
+Within the class of hierarchical self-organized systems, we investigate the interplay of structure and function, associated with the emergence of complex tree structures in disordered environments.
+Using an algorithm that creates and searches trees of real-world patterns, our work stands at the intersection of statistical physics, machine learning, and network theory.
+We resolve the network properties over multiple phase transitions and across a continuity of scales, using the von Neumann entropy, its generalized susceptibility, and the recent definition of thermodynamic-like quantities, such as work, heat, and efficiency.
+We show that scale-invariance, i.e.
+power-law Laplacian spectral density, is a key feature to construct trees capable of combining fast information flow and sufficiently rich internal representation of information, enabling the system to achieve its functional task efficiently.
+Moreover, the complexity of the environmental conditions the system has adapted to is encoded in the value of the exponent of the power-law spectral density, inherently related to the network spectral dimension, and directly influencing the traits of those functionally efficient networks.
+Thereby, we provide a novel metric to estimate the complexity of high-dimensional datasets.
+Trees, Networks, Thermodynamics, Statistical physics, Machine Learning, von Neumann entropy, Specific heat
+I Introduction
+Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully definite answer.
+In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of, on the one side, the optimal ability to store information about the environment in an accessible manner, on the other side, the ability to transmit information across the structure [2] (see Methods Sec.
+II.1.2).
+For instance, in living systems, information about the environment flows into the system’s genome through natural selection.
+The genome encodes such information into the organism’s structural traits, which in turn reflect the functional capability of the system.
+Thus, the system’s very own structure may be regarded as the internal representation of the environment [3, 4, 5, 6, 7] and its the structural traits could themselves enclose the complexity of the environmental conditions the system needs to adapt to.
 How these traits are optimized to maximize the fitness to the environment is an open question.
-The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines.
-Here we challenge this hypothesis.
-In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
-We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13].
-The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network.
-At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern.
-The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern.
-The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.
-Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$.
-For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out.
-For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created.
-The algorithm produces exclusively trees, i.e.
-loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively.
-Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
+The hypothesis of optimal functioning at the border of order and disorder suggests critical-like features as the potential candidates to obtain functional advantages [8, 9, 10, 11, 12, 13, 14] and maximize the fitness to the outer world, or, when in the context of learning machines, to obtain a better performance in a statistical generalization problem [10].
+Among the various attributes exhibited by truly critical systems and systems with critical-like properties, especially relevant to our work is the property of scale-invariance, exhibited through the power-law distribution of some meaningful quantity of the system under consideration.
 Figure 1: Tree structures generated by the Portegys algorithm.
@@
 Classes are indicated by the 10 different colors in the color bar.
-The use of a hierarchical topology, onto which a high-dimensional dataset is mapped, offers insights into the structured disorder of the data [15] on the one hand, while on a practical perspective, it limits the number of comparisons needed in matching an unknown pattern to a prototypical example.
-Namely, the algorithm introduced by Portegys is equipped with a pattern-searching method described in Section 3.1.
-Depending on the structural traits of the networks, an optimal trade-off between fast search of the best-matching pattern and accuracy in terms of node-class labels can be defined.
-The Portegys algorithm has the desirable properties of being simple, explainable, and subject to a single control parameter $r$.
-Thanks to these features, we can work on the bridge between three disciplines: statistical physics, machine learning, and network science, and we can show that information-theoretical tools such as the von Neumann entropy and its derivative reflect the functional capabilities of the network.
-Notably, we observe some consistency between the thermodynamic efficiency and pattern-matching efficiency that we try to explain in terms of average frequency content in the self-organized network structure.
-2 Theoretical framework
-2.1 Statistical physics of networks
+Hierarchical trees enable us to investigate the interplay between structure and function, associated with the emergence of complex structures in disordered environments, within a relevant class of systems.
+Trees are common to many real-world systems, from tree roots and branches, to neuronal dendrites [15, 16, 17, 18, 19] and biologically inspired artificial neural networks [20, 21, 22, 23, 24, 25], as well as technological networks for organizing and searching collections of high-dimensional patterns [26, 27, 28].
+We use the Portegys algorithm [27, 28] to produce hierarchical trees that self-organize given a collection of high-dimensional vectors.
+The algorithm recursively organizes (structural aspect) high-dimensional vectors randomly drawn from a dataset into trees (see in Methods Sec.
+II.2).
+Using the metaphor introduced by Jacob for natural selection [29], the Portegys tree-generation process acts as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
+The algorithm integrates a search phase (functional aspect) where it navigates the tree structure to achieve efficient pattern-matching.
+The algorithm is simple, explainable, and subject to a single control parameter, $r$.
+Still, it is capable of generating a large variety of branching structures (Fig.
+1) and multiple phase transitions (Sec.
+III.3).
+Other models can generate hierarchies.
+Such as the model in Ref.
+[30] where a single parameter controls the branching and produces structures in much analogy to the ones in Fig.
+1, or the Galton–Watson stochastic process originally proposed to model the extinction of family names  [31, 32].
+Nevertheless, the Portegys algorithm integrates the functional aspect given by the pattern-matching task and includes the disorder of the external dataset, two key attributes in our analyses.
+Thanks to these features, we can work at the intersection of three disciplines: statistical physics, machine learning, and network science.
+We can show that information-theoretical tools such as the von Neumann entropy [33] and its generalized susceptibility [34, 35] (see Methods), while monitoring the structural properties of the system, reflect the functional capabilities of the network.
+We show that emerging efficient structures exhibit a scale-invariant spectrum [36], i.e.
+power-law Laplacian spectral density.
+We reproduce this result over multiple datasets, ranked based on their complexity.
+Our findings suggest that a large range of availability in the power law spectral density provides the advantage of a large repertoire of network “frequencies”.
+These can be exploited to achieve the functional purpose of the system efficiently, akin to critical-like systems exhibiting scale-invariance in the distribution of a relevant quantity [37, 11, 38, 13, 14].
+Such behavior is observed consistently whether we take efficiency as gauged by the task-agnostic network thermodynamic efficiency [1], estimating the trade-off between the diversity of information pathways and fast information propagation (see Sec.
+II.1.2 in Methods), or the pattern-matching efficiency (Sec.
+II.2.2 in Methods), quantifying the trade-off between accuracy and number of visited nodes within the pattern-matching task.
+Furthermore, the exponent of the power-law spectral density, related to the network generalized dimension [39, 40, 41, 42], results to be driven by the complexity of the environment the system adapted to, directly influencing the traits of those functionally efficient networks.
+Finally, from a different perspective, using a hierarchical topology to map the distribution of high-dimensional vectors of a dataset provides insights into the structured disorder of the data [43].
+In particular, we measure the dataset network’s spectral dimension and show that it estimates the dataset’s complexity.
+Thus, among the other results, our work provides an alternative way to gauge the dataset’s complexity [43, 44, 45, 46, 47, 48].
+II Methods
+II.1 Theoretical framework
+II.1.1 Statistical physics of networks
 Information spreading over networked structures is governed by the Laplacian matrix [49] and its spectrum of eigenvalues encodes many relevant topological properties of the graph [50, 51].
@@
 Thus, for any state at $\tau$ it holds $|\phi(\tau)>=\sum_{i}^{N}e^{-\tau\lambda_{i}}|k_{i}><k_{i}||\phi(0)>$.
-This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum-Laplacian, while $\tau$ can be interpreted as the "wavelength" $l\sim\tau$ of the diffusion modes, thus it defines an operational resolution scale for observations [3].
+This relation remarks that the eigenvectors of the Laplacian are akin to the wave vectors in the Fourier space defined by the spectrum of the continuum Laplacian.
+At the same time, $\tau$ can be interpreted as the “wavelength”, $\ell\sim\tau$, of the diffusion modes, thus it defines an operational resolution scale for observations [35].
 Moreover, the ensemble of accessible diffusion modes at time $\tau$ can be constructed by defining a probability measure over the accessible modes decomposed by the spectral theorem.
@@
 The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location after time $\tau$ [54].
-Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
+Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity allowing its eigenvalues to be interpreted as probabilities.
 Finally, the von Neumann spectral entropy can be introduced [33] as a multi-scale descriptor of the network structure:
@@
 where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$.
-The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes.
+The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time, $\tau$, and is bounded between $[\log C_{conn},\leavevmode\nobreak\ \log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes.
 Thus the von Neumann entropy, $S$, as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [34].
-Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
-The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [22, 3].
-For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network.
-The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network.
-Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale.
+Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ and the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
+The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network.
+By the inspection of the entropy, $S$, and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [34, 35].
+For instance, a single large peak of the specific heat, $C$, identifies a single scale at which there is a fast and unique information diffusion (or information loss) across the network.
+The latter is the typical situation of a star graph, depicted in Fig.
+1, or an Erdös-Rényi network.
+Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the whole network, thus revealing a single trivial scale.
 On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate.
-Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3].
-The latter sentence makes for a definition for the scale-invariant networks [23].
-The specific heat plateau height, $C_{0}$, is also related to the spectral dimension of the network $d_{s}=2C_{0}$ [3, 23].
-From these considerations, a Laplacian-Renormalization group theory has been recently developed allowing for a coarse-graining of the network structures into super-nodes unraveling the mesoscale structures [3, 22].
-In this context, the peaks of the specific heat, $C$, or a marked change in its slope, identify the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [19].
-To conclude, we can intuitively use the diffusion parameter $\tau$ to resolve the network topological structures at different resolution scales.
-Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the "wavelengths" of its modes.
-2.2 Thermodynamics of networks
+Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed, $\mathcal{P}(\lambda)\sim\lambda^{\gamma}$, in a sufficiently extended scale-interval [35].
+The latter sentence makes for a definition for the scale-invariant networks [36], characterized by a scale-invariant information propagation across the network resolution scales.
+The the specific heat plateau height, $C_{0}$, is also related to the spectral dimension, $d_{s}=2(\gamma+1)$ [39, 41], according to the relation $d_{s}=2C_{0}$ [35, 36], which implies $C_{0}=\gamma+1$.
+The spectral dimension generalizes the concept of the Euclidean dimension to disordered structures and coincides with the Euclidean dimension for lattices.
+For instance, we expect $C_{0}=1/2$ for 1d-lattice; $C_{0}=1$ for the regular grid with periodic boundary conditions; and $C_{0}=4/6$ for random trees with finite degree variance $<k^{2}>$ showcasing fractal dimension $d_{s}=4/3$ [36].
+From these considerations, a Laplacian-Renormalization group theory has been recently developed, allowing for a coarse-graining of the network structures into super-nodes and for the unraveling of the mesoscale structures [35, 34].
+The theory relies on information diffusion to identify informationally equivalent “blocks” akin to the ones proposed by Kadanoff [55], and moves to the $k$-space to explain the procedure akin to Wilson’s formulation [56].
+In this context, the first peak of the specific heat, $C$, or a marked change in its slope, identifies the most convenient resolution scales where to perform the coarse-graining over the informationally connected nodes, as it allows the iteration of the procedure without an excessive network reduction, and, when the network is scale-invariant, maintaining the dilation symmetry across the renormalization group procedure [37].
+To conclude, we can intuitively use the diffusion parameter, $\tau$, to resolve the network topological structures at different resolution scales.
+Thus, the necessity of analyzing complex networked systems at all levels is solved by resorting to the conjugate Fourier space, the k-space, defined by the Laplacian spectrum and the “wavelengths” of its modes.
+II.1.2 Thermodynamics of networks
 The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables.
-To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1].
+In this regard, network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1].
 Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges diminishes the network’s entropy.
-Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits.
-The entropic cost is given by:
+Lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits.
+Thus, the entropic cost is given by:
 $Q=-\frac{\delta S}{\tau}=-\frac{\log N-S}{\tau}\leq 0$ (3)
@@
 In our work, we use this variational principle to explain the structural properties of the network subclass composed of heterogeneous hierarchical trees.
-An additional element is that these trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns.
-The specific heat is used here to identify which is the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime.
-For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=C_{0}/\tau$ [3, 23], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{C_{0}}{W}$.
-3 Methods
-Our results are evaluated over subsets of four different datasets: the NIST [25], the MNIST [26], the FashionMNIST [27], and the CIFAR10 [28] datasets.
+The specific heat is used to identify the resolution scale at which such a variational principle is effective for our subclass of networks during the transition to the scale-invariant regime.
+For scale-invariant networks, for some $\tau$-scales, $U=<\lambda>_{\tau}=C_{0}/\tau$ [35, 36], hence in this regime Equation 5 becomes $\eta=\frac{1}{\tau}\frac{C_{0}}{W}=\frac{1}{\tau}\frac{d_{s}}{2W}$.
+An additional element of this work is that the investigated trees are required to encode the diversity expressed in the environment they are in contact with, defined by a real-world dataset of patterns, $\mathcal{D}$.
+II.2 Portegys algorithm
+II.2.1 Network growth process
+To generate structures based on real-world disorder, we use the algorithm introduced by Portegys [27] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [28].
+The Portegys algorithm we use is the extended version with $k=5$, where the distance between patterns is the Euclidean distance.
+Given the external constraints, the single control parameter, $r$, and a high-dimensional vector collection, the dataset $\mathcal{D}$, the system seeks the best-fitness configuration to map the dataset’s disorder into a tree, i.e.
+a network that is loop-less and sparse, $|E|=N-1$, where $|E|$ is the number of links, and $N$ the number of nodes.
+The tree-formation protocol reorganizes raw data using a recursive method in which samples are randomly selected and inserted into the growing tree network.
+At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern.
+The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern.
+The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space.
+Figure 2: Search method based on relative distance.
+In (a) the growth process is illustrated.
+At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$.
+The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node.
+Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node, $P$, and the new pattern.
+In (b) the search method based on relative distance is illustrated.
+Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$.
+The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$.
+In the depicted case, patterns closer to the search pattern, $S$, are expected to be in the subtree of $C_{2}$.
+Adapted from [58].
+II.2.2 Search phase
+After the tree formation described in Sec.
+II.2.1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach.
+The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure.
+The growth process, described in Sec.
+II.2.1 and depicted in Fig.
+2(a), ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels.
+This provides the possibility of using a search technique based on relative distances.
+The search method is depicted in Fig.
+2(b) and is an iterative process that begins at the root-node pattern.
+Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtrees, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Fig.
+2(b).
+The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child, $C_{i}$, collectively indicated by the gray circle, $C^{\prime}_{i}$, in Fig.
+2(b).
+Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties.
+If $d(S,C_{i})_{search}$ is less than zero then it is set to zero.
+In Fig.
+2(b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
+II.3 Datasets
+Our results are evaluated over subsets of four different datasets ranked based on their complexity (Sec.
+III.1): the NIST [59], the MNIST [60], the FashionMNIST [61], and the CIFAR10 [62] datasets.
 These datasets contain randomly collected instances of image samples out of 10 classes.
-It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (’sub-style patterns’) such that the sample point cloud can be ordered naturally in a network structure.
-Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work.
-In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32].
-In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each.
-The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy.
-We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
+It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (“sub-style patterns”) such that the sample point cloud can be ordered naturally in a network structure.
 The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms.
@@
 In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently.
-Samples from the FashionMNIST and MNIST are always binarized, while the samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal.
-In the Supplementary we show also the same results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST.
-The 10 classes are equally represented both in the test and in the training set.
-The datasets used in our experiments exhibit different levels of complexity, see Figure 2.
-We estimate with multiple methods the dataset complexities on the same subset of 2000 training samples used in the rest of the manuscript.
-The first heuristic approach is based on the test accuracy of a small multilayer perceptron (2 layers with 128 units each).
+Samples from the FashionMNIST and MNIST are always binarized.
+The samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal.
+In the Supplemental Material [63], we show results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST.
+The 10 classes are equally represented in both the test and training set.
+III Results
+III.1 Dataset complexity
+The datasets used in our experiments can be ranked based on their estimated complexity, see Fig.
+3.
+We estimate each dataset’s complexity with multiple methods on the same subset of $2000$ training samples used in the rest of the manuscript.
+The first heuristic approach is based on the test accuracy of a small multilayer perception (MLP) (2 layers with 128 units each).
 In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes.
@@
 Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset.
-The second method we use is the computation of the Lempel-Ziv complexity [29] of the aggregated training set.
-In this case, samples need to be binarized.
-The third method concerns the measure of the intrinsic dimension [15, 32] of the dataset using the Maximum Likelihood Estimation approach introduced in [30] with the correction proposed in [31].
-A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the "world" of our problem, for instance, handwritten digits, it spans a lower dimensional manifold.
+The second method we use is the computation of the Lempel-Ziv complexity [45] of the aggregated training set. In this case, the samples need to be binarized.
+The third method concerns the measure of the intrinsic dimension [43, 48] of the dataset using the Maximum Likelihood Estimation approach introduced in Ref.
+[46] with the correction proposed in Ref.
+[47].
+A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the “world” of our problem, for instance, handwritten digits, it spans a lower dimensional manifold.
 Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [43, 48].
 We set the hyper-parameter $n=3$.
-While in Section 4 we discuss the full range of $r$, in Section 5 we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$.
-The Portegys algorithm we use is the extended version with $k=5$ introduced in [13] where the distance between patterns is the Euclidean distance.
-Figure 3: Search method based on relative distance.
-In (a) the growth process is illustrated.
-At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$.
-The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node.
-Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node $P$ and the new pattern.
-In (b) the search method based on relative distance is illustrated.
-Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$.
-The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$.
-In the depicted case, patterns closer to the search pattern $S$ are expected to be in the subtree of $C_{2}$.
-Adapted from [33].
-3.1 Pattern-matching efficiency
-After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach.
-The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure.
-The growth process, described in Section 1 and depicted in panel (a) of Figure 3, ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels.
-This provides the possibility of using a search technique based on relative distances.
-The search method is depicted in panel (b) of Figure 3 and is an iterative process that begins at the root-node pattern.
-Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtree, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Figure 3.
-The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child $C_{i}$, collectively indicated by the gray circle $C^{\prime}_{i}$ in panel (b).
-Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties.
-If $d(S,C_{i})_{search}$ is less than zero then it is set to zero.
-In panel (b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
-It is natural to define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
+Figure 3: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work.
+In green an estimate is given computing the Lempel-Ziv complexity [45], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [46, 47, 48].
+In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each.
+The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy.
+The CIFAR10 dataset is estimated as the most complex, while the NIST dataset is the least complex among the four.
+III.2 Pattern-matching efficiency
+Mapping a high-dimensional vector collection into a hierarchical tree structure limits the number of comparisons needed to match an unknown pattern to a prototypical example.
+Namely, the algorithm introduced by Portegys [27, 28] is equipped with a pattern-searching method illustrated in Fig.
+2 and described in Sec.
+II.2.2, representing the functional aspect of the system.
+We define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
 $\theta=\frac{\text{A}}{\log N_{dist}}$ (6)
 the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes.
-For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases, potentially decreasing the class-match accuracy.
-When defining the pattern-matching efficiency we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or terms of general resources.
-Hence, the efficiency $\theta$ quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in panel (c, f) of Figure 4 for the MNIST and CIFAR10 respectively.
-Note in panel (d) that the CIFAR10 dataset has a peak for accuracy which we interpret by saying that for complex datasets (see also the case of the FashionMNIST in Section S7) some network structures do possess the ability to better encode information of the "world" in their traits.
-We postpone the discussion of this finding to Sections 5 and 6.
-Figure 4: Pattern-matching efficiency in (c, f) measured over the NIST and CIFAR10 respectively.
-The efficiency $\theta$ is defined as the trade-off between the classification accuracy, (a, d), and the logarithm of the number of distances computed to find the best match, (b, e).
-The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency $\theta$.
+For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$; the number of computed distances decreases for increasing $r$, potentially lowering the class-match accuracy.
+Nevertheless, we highlight Fig.
+4(d) and note that the CIFAR10 dataset exhibits the global accuracy peak at an intermediate value of $r$.
+We interpret this observation by saying that for some datasets (see also the case of the FashionMNIST in Sec.
+S7 of the Supplemental Material [63]) specific network structures do possess the ability to better encode information of the “world” in their traits.
+We postpone the discussion of this finding to Secs.
+III and IV.
+When defining the pattern-matching efficiency, we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or in terms of general resources.
+Hence, the efficiency, $\theta$, quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in Fig.
+4(c),(f) for the MNIST and CIFAR10 respectively.
+While in Sec.
+III.3 we discuss the full range of $r$.
+In the rest of the manuscript, we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover, the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$ and $N=2000$.
+Figure 4: Pattern-matching efficiency in (c),(f) measured over the NIST and CIFAR10 respectively.
+The efficiency, $\theta$, is defined as the trade-off between the classification accuracy, (a),(d), and the logarithm of the number of distances computed to find the best match, (b),(e).
+The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency, $\theta$.
 The shaded grey area designates the standard deviation.
-The purple shaded area in panel (c, f) indicates a tolerance $r$-domain of optimal efficiency.
+The purple shaded area in panel (c),(f) indicates a tolerance $r$-domain of optimal efficiency.
 The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area.
@@
 The 10-digit classes are equally represented both in the train- and test sets.
-4 Phase portrait
-Figure 5: Multiple phase transitions in hierarchical tree networks.
-In (a) the specific heat, $C$, across the range $r\in[0,2]$.
-At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat.
-The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line.
-For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot C_{0}$, where $C_{0}$ is the plateau height in (c).
-Note that at $r\sim 0.8$ the spectral dimension is $d_{s}=4/3$ as expected for random trees with finite variance in the node degree.
-In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.35,~{}0.40,~{}0.70,~{}0.80,~{}0.90,2.00\}$.
-Vertical blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively.
-In (d), the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate.
-In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right.
-All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations.
-Figure 6: Order parameters.
-In (a), the modularity, $Q$, resolves the transition involving the mesoscale structure.
-In (b), the normalized diameter, $\hat{d}=d/m$ resolves the large-scale transition.
-Their respective variances (dashed lines) identify the transition points.
-In (c), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-domain, i.e.
-the number of nodes included in the scale-invariant region.
-Note the discontinuity at the transition from star to amorphous.
-Curves in (a-b) are realized for networks of $N=500,~{}1000$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
-Curves in (c) are for networks of $N=500$.
+III.3 Phase portrait
 […]
-The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$.
-Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in (a) and (c) of Figure 5.
+[…]
+The Portegys algorithm produces three prototypical structures as those depicted in Fig. 1 and Fig. 5(b), depending on the value of the control parameter, $r$ (see in Methods Sec. II.2).
+For $r\rightarrow 0$, every node is linked to the root node and a star graph turns out (see in Fig.
+1 the central node, which is the first node selected in the growth process).
+For some intermediate value of $r$, a ragged tree structure with progressively smaller distances between the patterns is produced; while for $r\rightarrow 2$, a 1d-lattice is created.
+[…]
+The second change of phase is the transition from the star graph to the amorphous configuration at $r\simeq 0.35$.
+Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in Fig.
+5(a),(c).
 For $r\rightarrow 2$, the plateau would include the full 1d-lattice in the thermodynamic limit $N\rightarrow\infty$.
@@
 After the phase transition, a second peak of the specific heat emerges at larger $\tau$ scales.
-This second peak of $C$ in Figure 5.(c) identifies the scale at which the information is integrated between the distinct modules.
-In other words, the second peak reflects the "degree" of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes.
+This second peak of $C$ in Fig.
+5(c) identifies the scale at which the information is integrated between the distinct modules.
+In other words, the second peak reflects the “degree” of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes.
 Note that we exclude both the clusters, i.e.
-triangles, and higher-order cycles from our discussion as the tree networks we address are purely loop-less.
-This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies always occur in the scale-invariant region of the phase plane, and in some cases, they occur in the vicinity of this transition border $r\sim 0.35$.
+triangles, and higher-order cycles from our discussion, as the tree networks we address are purely loop-less.
+The star graph and amorphous network phase space regions are those investigated in this work.
+Because the peaks in the thermodynamic and pattern-matching efficiencies always occur in the amorphous and scale-invariant region of the phase plane.
+In some cases, they appear in the vicinity of the transition border between the two at $r\simeq 0.35$.
 This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network.
-Indeed, we observe in Figure 6.(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition.
+Indeed, we observe in Fig.
+6(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition.
 Nevertheless, a phase transition exists and an appropriate order parameter needs to be employed to identify it.
 This phase change occurs at smaller resolution scales with respect to the first phase transition we discussed.
-Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation in the specific heat in Figure 5.(a),(c)).
-Thus, we can identify this transition using as an order parameter the modularity [35], $Q$, of the tree which we show in panel (a) of Figure 6.
-We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36].
+Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation of the specific heat in Fig.
+5(a),(c)).
+Thus, we can identify this transition using as an order parameter the modularity [65], $Q$, of the tree which we show in Fig.
+6(a).
+We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in Ref.
+[66].
 We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor.
-This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
-As a second possible order parameter, for the mesoscale phase-transition, we show in Figure 6.(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks.
+Thus, a mesoscale resolution is required to monitor the structural changes occurring across the star graph phase to the amorphous scale-invariant phase transition.
+This motivates the choice of the mesoscale $\tau\sim 10$ employed in Sec.
+III.
+As a second possible order parameter, for the mesoscale phase-transition, we show in Fig.
+6(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ specific heat peaks.
 This order parameter counts the normalized (log)number of informationally connected nodes across the $\tau$-scales of the specific heat plateau.
@@
 We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
-As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see Figure 5.(d)).
+As already discussed in Sec.
+II.1.1, also along the $\tau$-direction an informational phase transition exists [34, 35] delimited by the peaks of $C$ at $\tau^{(1,2)}_{peak}$.
+Using the specific heat peaks, we can distinguish three dynamical phases (see Fig.
+5(d)).
 For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase.
 Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [35, 36].
-5 Results
-Figure 7 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
-Panels (a)-(d) depict the thermodynamic efficiency, $\eta$, and panels (e)-(h) depict the pattern-matching efficiency, $\theta$.
-The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in (e)-(h) indicate a $r$-domain tolerance for $\theta$.
-The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
-Note the consistency in the global maxima of $\theta$ and $\eta$ in terms of optimal $r$ for the NIST and the FashionMNIST.
-See the $r$-range spanned by the purple-shaded areas in panels (e)-(h) and the vertical dashed lines in panels (a)-(d).
-This consistency in the maxima suggests that the two efficiencies might be related, for the specific context of this algorithm.
-In the Section 6 we discuss this matter.
-The topological complexity of the networks increases at a critical $r$ from the trivial traits of a star graph to the fractal structure of the amorphous network.
-Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset.
-Such transition is also depicted in Figure 8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 7.
-See also Section S4 which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
-We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 7).
-This is an intrinsic characteristic of the dataset, as it is only mitigated for larger networks, as shown in Section S2 of the Supplementary.
-This intermediate peak is sufficiently strong to be reflected on a local maximum of $\eta$, in Figure 7.(c).
-Although this local maximum in $\eta$ is not detected by $\theta$ in Figure 7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area when we compare $\eta$ and $\theta$ peaks in terms of $r$.
-This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Figure 7.(c).
-Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Sections 4 and 6.
-Such critical resolution scale is identified by the bifurcation of the specific heat peaks in panels (i)-(l) of Figure 7 and in panels (a)-(d) of Figure 8.
-In the seminal work [1], it is suggested to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
-the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales.
-In fact, this eigenvalue should provide the time scale where the diffusion process has reached equilibrium.
-Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$.
-Panels (a)-(d) of Figure 8 depict the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves.
-The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat.
-It is crucial to identify the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in Figure 5.(b).
-We consider such resolution scale of significant importance for the computation of the efficiency $\eta$ across the transition as it defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3].
-In Section S3 of the Supplementary Material the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e.
-the red curves depicted in panels (a).
-Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
-Finally, note in Figure 8 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$.
-We remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only, and the latter shapes the degree distribution only, as the number of nodes and edges is fixed across the transition.
-We note that the Portegy’s algorithm stays in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
 Figure 7: Thermodynamic & pattern-matching efficiency.
-In (a)-(d) the thermodynamic efficiency, $\eta$, as a function of $r$.
-The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l).
-The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant "length" scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure.
-In (e)-(h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure).
+In (a–d) the thermodynamic efficiency, $\eta$, as a function of $r$.
+The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i–l).
+The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant “length” scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure.
+In (e–h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure).
 The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line.
-In panels (b)-(f) and (d)-(h) a noticeable cohesion of the peaks of the two efficiencies is observed.
+In panels (b),(f) and (d),(h) a noticeable cohesion of the peaks of the two efficiencies is observed.
 Vertical dashed lines indicate the global maxima of the efficiencies.
 The efficiency peaks, both $\eta$ and $\theta$, are always in the amorphous region of the phase plane.
-For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star-graph to the amorphous region.
+For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star graph from the amorphous region.
 This agrees with the argument that a fairly simple dataset does not benefit from the vicinity of the phase transition.
 Quantities are averaged over 50 independent simulations.
+III.4 Efficiency
+In Fig. 7, panels (a–d) depict the thermodynamic efficiency, $\eta$, and panels (e–h) depict the pattern-matching efficiency, $\theta$. See Equations 5 and 6 respectively.
+The thermodynamic efficiency [1], $\eta=(Q+W)/W$, quantifies the trade-off between two quantities.
+The former, $Q=-\delta S/\tau$, is proportional to the variation of entropy content, $\delta S$, from the case of isolated nodes (see Methods).
+Thus, it estimates the ability to express diversity to a stochastic perturbation, or in other terms, the internal representation capabilities provided by the network topological traits.
+The latter is the variation of information flow with respect to a network of isolated nodes, $W=\delta F$; so to speak, how fast information travels across the network (see Methods).
+The pattern-matching efficiency, $\theta$, defined in Equation 6, quantifies the trade-off between accuracy and number of visited nodes in the pattern-recognition task, representing the functional aspect of the algorithm (see Methods).
+The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in panels (e–h) indicate a $r$-domain tolerance for $\theta$.
+The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
+The two efficiencies $\theta$ and $\eta$ show some consistency in the x-range of highly efficient structures, see panels (b),(d) related to the NIST and the FashionMNIST.
+If we include points in the tolerance $r$-range spanned by the purple-shaded areas in panels (e–h), we observe the same consistency for all datasets when $\eta$ is computed at $\tau=10$ (red dashed lines in panels (a–d)).
+This consistency in the maxima suggests that the two efficiencies might be related, in the specific context of the algorithm here investigated.
+In the Sec.
+IV we discuss this matter.
+The specific resolution scale, $\tau\sim 10$, at which we compute $\eta$ is the scale involved in the mesoscale structural transition from star graph to amorphous scale-invariant, that we discussed in Sec.
+III.3 and Fig.
+6.
+Such critical resolution scale is identified by the bifurcation of the specific heat peaks in Fig.
+7(i–l) and in Fig.
+8(a–d).
+In the seminal work in Ref.
+[1], it is suggested to use as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e.
+the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales where to compute the thermodynamic efficiency, $\eta$.
+The Fiedler eigenvalue provides the largest network time scale where the diffusion process has reached equilibrium.
+Nevertheless, $\tau_{diff}$ seems to only approximate the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$, especially in the star-graph region for small $r$.
+Panels (a–d) of Fig.
+8 depict the $\tau_{diff}$ curves in red, next to the $\tau^{(1,2)}_{peaks}$ curves.
+The key difference is the thermal ensemble average: a weighted average, that employs all the eigenvalues to compute the specific heat and identify its peaks.
+Differently from what was proposed in Ref.
+[1], we don’t use $\tau_{diff}$, or $\tau_{peak}^{(2)}$ in Fig.
+8.
+Because that would mean adopting the largest possible resolution, which would not monitor the structural mesoscale changes involved in the transition, from the single trivial scale of the star graph to the more complex topological traits of the amorphous networks in the scale-invariant regime.
+The resolution scale, $\tau\sim 10$, is shared across all topologies in the phase plane and, remarkably, defines an “ultra-violet” cut-off in the characteristic frequency of the networks across the transformation from one phase to the other.
+This is in analogy to the process of integrating out the large eigenvalues responsible for the finer details, as discussed in the context of the Laplacian-Renormalization group [35].
+Note in Fig.
+S1 of the Supplemental Material [63] that our choice of $\tau=1/\lambda\sim 10$ is supported by the exclusion of the resolution region where the power-law behavior of the spectral density, $\mathcal{P}(\lambda)$, is lost.
+In Sec.
+S3 of the Supplemental Material [63], the efficiency, $\eta$, is computed in the range interval $\tau<10$.
+Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching efficiency.
+Moreover, we observe that, while the value of $r$ where a structural transition appears is dependent on the dataset, i.e.
+$r=0.40$ for the NIST, $r\simeq 0.40$ for the MNIST, $r\simeq 0.35$ for the FashionMNIST, and $r\simeq 0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is the same for all the investigated datasets.
+Finally, we note that the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplemental Material [63] Sec.
+S2).
 Figure 8: Role of the ensemble average on the estimation of the $\tau$-scale.
-Panels (a)-(d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$.
-$\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 7.
+Panels (a–d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$.
+$\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Fig.
+7.
 $\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$.
-Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i)-(l) of Figure 7.
+Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i–l) of Fig.
+7.
 At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$.
@@
 The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
-5.1 Performance at the emergence of scale-invariance
-In this section, we discuss the relation between the location of the pattern matching efficiency, $\theta$, on the phase plane of the algorithm and the complexity of the dataset, see Figure 2.
-We turn our attention to the panels (e)-(h) and (i)-(l) in Figure 7 and panels (a)-(d) in Figure 8.
-For the CIFAR10 and FashionMNIST datasets, there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the peaks of each of the two $\theta$ efficiencies are observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively.
-Note that for the FashionMNIST, also the thermodynamic $\eta$ peaks in the same range of $r$.
-On the other hand, for the MNIST and NIST datasets, $\theta$ peaks in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$.
-The term "far away" is quantified by the purple-shaded areas in panels (e)-(h) of Figure 7.
-Note the correspondence of the peaks of $\eta$ and $\theta$ for the two datasets in terms of $r$.
-The latter observation on the correspondence of $\theta$ and $\eta$ also holds for the MNIST, if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
-To explain the varying locations of peak efficiencies, whether near the transition border or farther from it, we refer to an argument in the literature [37, 10, 38].
-A system needs tuning, or self-tuning, at the edge of a phase transition only when the task demands a high level of complexity, specifically, this was investigated in the context of second-order phase transitions [37, 10].
-For the more complex datasets CIFAR10 and FashionMNIST, the optimal $\theta$ values are closer to the transition border.
-In contrast, the optimal values for the simpler NIST and MNIST datasets are located farther from their respective transition points.
-This observation also holds for the thermodynamic $\eta$ for the FashionMNIST, MNIST, and NIST.
-Moreover, the MNIST, an intermediate dataset in terms of complexity, is an interesting case.
-It shows two maxima for the thermodynamic efficiency, $\eta$.
-The global maximum in Figure 7(c) is far away from the critical value of $r$ illustrated by the bifurcation in panel (c) of Figure 8, while the local maximum is at $r\sim 0.35-0.40$, which corresponds to the critical value of $r$ indicated by the bifurcation.
-These observations align with the hypothesis that the system exploits its proximity to the phase transition boundary only when presented with a relatively complex task.
-The overall discussion highlights the importance of the dataset under analysis, particularly its disordered distribution of points within a high-dimensional feature space.
-Finally, we observe that while the value of $r$ where a structural transition appears is dependent on the dataset, i.e.
-$r=0.50$ for the NIST, $r=0.40$ for the MNIST, $r=0.35$ for the FashionMNIST and $r=0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is roughly the same for all the four datasets.
-Interestingly, the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary Section S2).
-We speculate about the possibility that this is due to the constraint on sparsity as it is an invariant feature over the cases investigated in this work.
-6 Discussion
+III.5 Datasets’ phase plane
+The topological complexity of the networks increases at a critical $r$, from exhibiting the trivial traits of a star graph to the structure of the amorphous network with a roughly power-law spectrum.
+Thus, in the phase plane, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau between the first and the last peak.
+This is illustrated in Fig.
+7(i–l) for each investigated dataset.
+Such transition is also depicted in Fig.
+8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in Fig.
+7(i–l).
+See also Sec.
+S4 in the Supplemental Material [63] which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
+The MNIST, an intermediate dataset in terms of complexity, is an interesting case.
+We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Fig.
+7(k)).
+This is an intrinsic characteristic of the dataset, as this feature is only mitigated for larger networks, as shown in Figs.
+S2 and S7 of the Supplemental Material [63].
+This intermediate specific heat peak is reflected on two maxima for the thermodynamic efficiency, $\eta$, and a two-step transition in the amorphous region, in Fig.
+7(c),(k).
+The local maximum is at $r\in[0.35-0.45]$, depending on the specific $\tau$-scale.
+Such a range aligns with the transition at $r\simeq 0.40$ indicated by the bifurcation of the two specific heat peaks in Fig.
+8.
+The global maximum in Fig.
+7(c) is close to the second-step transition at $r\simeq 0.6$.
+Although the local maximum in $\eta$ is not detected by $\theta$ in Fig.
+7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area, used to compare $\eta$ and $\theta$ peaks in terms of $r$.
+This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Fig.
+7(c).
+Finally, note in Fig.
+8 the slow down of diffusion due to the increasing “volume” of the network discussed in Sec.
+III.3 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$.
+We remark that $\tau^{(2)}_{peak}$ is a function of solely the parameter $r$.
+The latter shapes solely the degree distribution, as the number of nodes and edges is fixed across the phase plane and fixed in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
+III.6 Efficient performance at the emergence of scale-invariance
+In this and the next section, we discuss the relation between the location of the maximum pattern-matching efficiency on the algorithm’s phase plane (Fig.
+7(e–h)) and the dataset’s complexity (Fig.
+3).
+All optimal networks are found in the amorphous region of the phase plane where the dilation symmetry is observed in the spectral density and reflected in the specific heat plateau.
+To discuss this observation we refer to the following argument [67, 11, 68]: a system needs tuning, or self-tuning, at the edge of a phase transition when the task demands a high level of complexity.
+Specifically, this was investigated in the context of second-order phase transitions [67, 11].
+But what is the advantage of being close to a second-order phase transition?
+One possible advantage comes from the power-law behavior of the density function of some quantities, e.g.
+density of cluster sizes in percolation, or density of avalanche sizes and duration in systems with temporal dynamics.
+The power law provides the system with a large range of availability from the density of the quantity of interest, and the system may exploit such availability as a resource in the task, whether it is adaptation to a complex environment or other goal-oriented ones.
+In our case, the amorphous region of the phase plane, where the network traits display self-similar structures in a limited $\tau$-resolution range, exhibits the power law of the spectral density, governing the network frequencies.
+Thus, better performances in terms of efficiency, particularly pattern-matching efficiency, can be explained in terms of network frequencies, which provide effective resources for the real-world disorder structuring task.
+III.7 Spectral dimension estimates the dataset complexity
+We now examine the location of the optimal $\theta$ efficiency in the amorphous region of the phase plane for each dataset.
+Thus, for each dataset, we inspect the traits of the highly efficient structures.
+We resort to the spectral dimension [39, 40, 41], $d_{s}$, which generalizes the concept of Euclidean dimension to disordered structures.
+It takes non-integer values between 1 and 3 and the same value of the Euclidean dimension in the case of d-dimensional lattices.
+The spectral dimension gauges the effective large-scale connectedness of disordered structures, with high values of $d_{s}$ indicating high topological connectedness, and is related to the asymptotic behavior of diffusion [40], but also to electrical transport and thermal stability of folded proteins [42]
+Depending on where the maximal $\theta$-efficiency is observed in the amorphous region of the phase plane in Fig.
+7, we see networks changing the value of the specific heat plateau, reflecting the spectral dimension, $d_{s}=2C_{0}$ [34], as we discussed in Sec.
+II.
+We also observe that the specific heat plateau is an approximately decreasing function of $r$ in the scale-invariant region of the phase plane in Fig.
+5(a).
+See also Sec.
+S4 in Supplemental Material [63].
+For complex datasets, such as the CIFAR10 and FashionMNIST, the optimal $\theta$s are closer to the transition border delimiting the star graph from the amorphous phase.
+In contrast, for the simpler NIST and MNIST datasets, the optimal values are farther from their respective transition points.
+This observation also holds for the thermodynamic efficiency, $\eta$, noted the correspondence of the peaks of $\eta(\tau=10)$ and $\theta$ for these datasets in terms of $r$.
+A special case is again the MNIST, with its two-step transition in the amorphous region, where local maximum $\eta$ are at the first transition, $r\simeq 0.35$, and the global maxima of $\theta$ and $\eta$ are at $r\simeq 0.6$.
+We measure the spectral dimension for each dataset (N=2000) by averaging over the exponents of the spectral density $\mathcal{P}(\lambda)\sim\lambda^{\gamma}$.
+The exponents are obtained fitting the power law in the range $\lambda\in[1\times 10^{-3},2\times 10^{-1}]$ for all those networks which better adapted to the dataset, i.e.
+those included in the purple $r$-range of maximal $\theta$-efficiency in Fig.
+7(e-h).
+The spectral dimension, $d_{s}=2(\gamma+1)$ [41], is plotted with red dots for each dataset in Fig.
+9.
+We removed the value $\gamma(r=0.35)$ when measuring the spectral dimension of the CIFAR10 because the spectral distribution was too close to the transition to obtain a good fit of the power law.
+We also estimated the spectral dimension by measuring the specific heat plateau [35, 36], $C_{0}$, over the resolution range $\tau\in[1,2]\times 10^{1}$.
+The measurements, $d_{s}=2C_{0}$, are shown with blue triangles in Fig.
+9.
+Fig.
+9 shows that the amorphous networks’ spectral dimension is a good estimator of the dataset complexity in Fig.
+3.
+For higher-complexity datasets (ranked from left to right, from the low complexity NIST to the high complexity CIFAR10) we have increasing spectral dimension, $d_{s}$.
+In other words, the spectral dimension provides a way to gauge the complexity of the dataset after applying the tree generation protocol.
+Moreover, we can assess that high-complexity datasets are arranged in topological structures that are more compact when compared to those of lower complexity.
+Figure 9: Datasets’ spectral dimension.
+The amorphous networks’ spectral dimension is a good estimator of the dataset complexity (Fig.
+3).
+The red points indicate the spectral dimension [41], $d_{s}$, measured from the power-law exponent of the spectral density.
+The blue dots are obtained by monitoring the specific heat plateau [35, 36], $d_{s}=2C_{0}$, averaged over the resolution $\tau\in[1,2]\times 10^{1}$.
+All points are averaged over the $r$-range of the phase plane of maximal $\theta$-efficiency, see the shaded purple areas in Fig.
+7(e–h), and over 50 independent simulations.
+IV Discussion
 The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [69].
-In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset - they are posed in contact with.
-Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed way as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine to achieve statistical generalization [41].
-When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42].
-Moreover, the same system needs to cope with the cost of sending information across its structure which can cause an enhanced exploitation of resources.
-In the context of machine learning, the cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g.
-class-matching.
-The recently developed theoretical tools of statistical physics for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy.
-As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
-In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure.
-The tree formation undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
-Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different levels of complexity.
-The optimal pattern-matching efficiency was always found in the amorphous regime, Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star-graph and the scale-invariant regimes.
-Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range.
-Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an ergodic process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
-In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$.
+In other words, systems, whether living systems or machine(-learning) systems, need to learn an internal representation of the complex world - or the dataset - they are posed in contact with.
+Ideally, to be competitive, such internal representation should be efficient in terms of desired function and available resources.
+It can be encoded in terms of abstract probability distributions [10, 11] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [13].
+Or in a mixed way, as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine able to achieve statistical generalization [70].
+When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli, possibly drawn from the same distribution they adapted to [71].
+Moreover, the same system needs to cope with the cost of sending information across its structure, which can lead to an enhanced exploitation of resources.
+In machine learning, the cost can be quantified by the number of computations required before the satisfying answer to a task, such as class matching, can be reported.
+Here, we used an algorithm [27] that creates and navigates trees of real-world patterns to investigate the intertwined role of structure and function in hierarchical self-organized tree systems in a simple setting.
+The algorithm, akin to living systems, adapts to the diversity of the environment it is in contact with, represented here by the real-world dataset.
+It self-organizes structuring the disordered collection of high dimensional vectors into its network traits.
+The tree formation process undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
+Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets exhibiting distinct complexity levels.
+The recently developed theoretical tools of statistical physics for networks [33, 54, 72, 34, 35, 1] permit to quantify the diversity expressed in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy.
+As the speed of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two.
+Thus, a variational principle gauged by the thermodynamic efficiency can be devised to explain why a certain structure emerges in place of another [1].
+In our work, this variational principle and the other tools provided by the theory are computationally employed to monitor the structural properties of sparse heterogeneous hierarchical trees produced by the Portegys protocol.
+Empirical evidence suggests that the thermodynamic and the pattern-matching efficiencies show some consistency, namely when we look at the global maxima of the two in the $r$-range.
+These two quantities involve two distinct processes that possess significant differences.
+On the one hand, the Portegys search phase is a directed process starting from the root node that navigates the structure under a homophily principle.
+On the other hand, the diffusion propagator describes an undirected ergodic process propagating from any possible node.
+Moreover, while the thermodynamic efficiency is defined based on general statistical physics arguments, and thus it is task-agnostic; the pattern-matching efficiency is defined based on the specific task under consideration.
+Still, we think it is reasonable to expect a similar behavior of the two efficiencies.
+Both quantities estimate a trade-off between the flow of a process and the encoding capabilities of the system’s topological traits.
+In both cases, the “flow” is quantified by a monotonic decreasing quantity as a function of $r$.
 For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$.
-It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4].
-They are not exactly alike as only the latter is an ergodic process; still, they can be used as estimates of the "flow".
+It is easy to see that these two quantities describe the flow in the network.
+The former counts the number of visited nodes across the traversing of the network by the Portegys search process.
+The latter is minus the logarithm of $Z$, with $Z\sim\mathcal{R}(\tau)$ where $\mathcal{R}(\tau)$ is the average return probability after time $\tau$ of a continuous time edge-centric random walker [54, 73].
+Thus even though they are not alike, they can still be used as proxies of the “flow”.
 We have discussed the denominator of the two efficiencies: $\eta$ and $\theta$.
 We now discuss the numerator.
-For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
-The matter is what makes a certain structure better suited in terms of accuracy to a given task, or equivalently in terms of a complex input-output mapping.
-Analogously to artificial neural networks, enlarging the number of parameters eventually results in achieving more complicated functions, even though this doesn’t imply a one-to-one relation between the two.
-In the case of a network structure, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e.
+For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se.
+The accuracy is not necessarily a monotonic decreasing function of the control parameter, $r$.
+It can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.
+The matter is what makes a certain structure better suited in terms of an objective, such as accuracy or efficiency, to a given task.
+Equivalently and more generally said, in terms of a complex input-output mapping.
+When dealing with artificial neural networks, we can either enlarge the number of parameters or adjust their values to achieve more complicated functions, even though neither implies a one-to-one relationship with better generalization performance.
+In the case of our network structures, the parameters are the network traits themselves, which in the Fourier space become the “network frequencies”, $\lambda_{i}$, i.e.
 the Laplacian eigenvalues.
-We can measure an effective number of parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5.
+We can measure a weighted average of the parameters by computing the thermal average of the frequencies contained in the network, precisely the numerator in Equation 5.
 The internal energy, $U$, is the Gibbs average of the Laplacian eigenvalues of the network at the operational resolution scale, $\tau$, tantamount to the inverse temperature, $\beta$, in statistical physics.
-Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau$ [3].
-The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, but we can let the structure self-organize given the value of the external control parameter, $r$.
-Notably, we have computed the thermodynamic efficiency at a specific resolution scale, $\tau$.
-We have used the specific heat $C=-dS/d\log\tau$ to identify the $\tau$-resolution which resolves the transition from the trivial structure, that is a single specific heat peak, to the emergence of more complex structures, where the specific heat shows some plateau reflecting the growth of a scale-invariant region where information is integrated at a constant rate.
-Such $\tau$-resolution scale is akin to an "ultra-violet" cut-off delimiting the relevant frequencies of the network structure across the transition to scale-invariance, i.e.
-from star graph to amorphous.
-The operational $\tau$-scale employed roughly identifies the scale where modules emerge in the network.
-For this reason, we could gauge the phase transition employing the network modularity, $Q$, as an order parameter.
-While the transition to scale-invariant permits the identification of the network resolution scale, $\tau$, where the thermodynamic efficiency needs to be measured to observe a peak, the corresponding transition value of the control parameter, $r$, does not always coincide with the global maximum in the efficiency.
-We explain this with the statement that a system benefits from the vicinity to a phase transition only when it confronts environments of higher complexity, while it does not benefit from it when the environmental complexity is lower.
-This is elucidated by the use of four different datasets of increasing complexity and supported by other works in the literature [37, 10, 38].
-In terms of practical utility, the methods presented could in principle provide diagnostics to tree-based approaches to classification such as hierarchical clustering, even though a limitation of this study is the investigation of only the Portegys algorithm.
-Given a pre-trained tree network based upon an unknown dataset, it might be possible to qualitatively identify the complexity of the dataset and the ability of the tree structure to cope with that complexity based on the position of the peaks of the pattern-matching and thermodynamic efficiency on the phase plane without performing any test phase.
-As discussed in this manuscript, if the trained network is “struggling” in the representation of the dataset, the peak of the pattern-matching efficiency might be at the phase transition boundary providing higher accuracy at the cost of a slower pattern search.
-Interestingly, our analysis of hierarchical tree networks could be applied, if further developed, to other topics in artificial intelligence.
-We find it particularly suited to the context of dendritic learning [44], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for biologically inspired computing.
-In the context of pure neuroscience, it would also be interesting to design experiments where the dendritic structure is related to the capability of the neuron to deal with complicated input-output mapping of signals.
+Thus, the Gibbs average frequency content is quantified by the internal energy, $U_{\tau}=\operatorname{Tr}[\hat{\rho}\hat{L}]=<\lambda>_{\tau}$, where for scale-invariant networks $<\lambda>_{\tau}=C_{0}/\tau=\frac{1}{2}\frac{d_{s}}{\tau}$ [35] in a limited $\tau$ range due to the finite size of network and the frequency cut-off that we discuss in the following.
+Thus, $U$ is related to the specific heat plateau, $C_{0}$, and, importantly, to the spectral dimension [41], $d_{s}$, which governs the spectral density.
+The difference to the case of the artificial neural network is that we can’t directly increase the number of parameters, given a fixed number of samples, or directly adjust the parameter values.
+Still, we can govern the distribution of frequencies with the control parameter, $r$, and let the structure self-organize.
+We have computed the thermodynamic efficiency at a specific resolution scale, $\tau=10$, identified by inspecting the specific heat, $C=-dS/d\log\tau$.
+At this $\tau$-scale, we could monitor the transition from the trivial structure, a single specific heat peak, to the emergence of complex amorphous structures, where the specific heat shows a plateau.
+The plateau reflects the growth of an approximately scale-invariant regime where information is integrated at a constant rate.
+Such $\tau$-resolution scale is akin to an “ultra-violet” cut-off, delimiting from above the relevant frequencies of the networks across the transition to scale-invariance.
+At the same time, it scans the scale where modules emerge in the (loop-less) network.
+For this reason, we could gauge the phase transition to the amorphous regime with the network modularity, $Q$, as an order parameter.
+Finally, this resolution $\tau$-scale is independent of the control parameter, the dataset, and the doubling of the number of nodes (see Supplemental Material [63]).
+Highly efficient networks were always found in the amorphous regime.
+From a functional point of view, we mention that several works in the literature [67, 11, 68] suggest that a system benefits from the vicinity to a phase transition critical point when it confronts environments of high complexity, while it does not benefit from it when the environmental complexity is lower.
+This partly aligns with our findings on the four datasets of increasing complexity.
+One key aspect of being close to a critical point is to take advantage of the functional benefits that a power law density related to some quantity of interest has to offer.
+In our case, for all datasets, we saw higher performance in terms of pattern-matching efficiency in the amorphous region of the phase plane, characterized by a power law spectral density.
+Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star graph and the scale-invariant regime.
+On the contrary, lower complexity datasets find their highly efficient structures on the opposite side of the amorphous phase plane region.
+Then what is the reason for a network embedding high dimensional patterns to be in one or the other location of the phase plane?
+By measuring the optimally efficient networks’ spectral dimension [39, 41], $d_{s}$, almost continuously changing in the amorphous regime, we could relate the dimensionality of the network to the complexity of the dataset it structured.
+We observed higher spectral dimension, reflecting large-scale connectedness, for networks mapping high-complexity datasets.
+On the contrary, easier datasets were mapped into more complex branching structures with a lower spectral dimension.
+Thus, importantly, we provide an alternative way to measure the complexity of a dataset: by mapping it into a tree network and then measuring its spectral dimension.
+Returning to the starting question of this manuscript—why a certain structure emerges in place of another—the answer suggested by our findings is: to manage the complexity of the environment it is adapting to.
+The generalization of this possible answer to systems other than the Portegys algorithm remains purely speculative and beyond the scope of this work.
+Still, the methods presented can in principle provide diagnostics to other tree-based approaches to classification such as hierarchical clustering.
+Suppose we are given a pre-trained tree network mapping an unknown dataset, we can estimate the complexity of the dataset that was used for training, along with the ability of the tree structure to cope with that complexity without performing any test phase.
+Suppose the trained network “struggles” in the representation of the dataset.
+In such a case, the spectral dimension is expected to be high, and the network is expected to exhibit higher connectedness.
+On the contrary, a network mapping a low-complexity dataset should exhibit branched structures and low spectral dimension.
+In the reverse direction, knowing the complexity of a dataset, we could initialize the search for the optimal control parameter, $r$, closer to values typically presenting one of the two transitions exhibited by the algorithm, reducing the number of simulations required to find the most efficient network structure.
+Interestingly, developments of our analysis and methods on hierarchical tree networks could be applied to other topics in artificial intelligence, neuroscience, and neuromorphic computing.
+We find it particularly suited to the context of biologically inspired [16, 17, 15] dendritic learning [20, 21, 22, 23], where input units are connected to output units via loop-less paths, to guide the design of the tree-like dendritic structures for parameter-efficient learning in artificial neural network [20, 21, 22, 23].
+Neurons can be modeled as the amorphous tree structures discussed in this work.
+It would be interesting to experimentally relate the dendritic structure, specifically its spectral dimension, to the capability of the neuron to deal with environmental stimuli and complicated input-output mapping of signals [15, 16, 17, 18, 19], or to produce signals with richer diversity [74, 71, 75].
+For instance, a neuron capable of dealing with the complexity of the stimulus it is subject to would be expected to exhibit traits with low spectral dimensionality and complex branching.
 Moreover, potential applications include the topic of neuromorphic computing [44, 76, 77] currently looking for a possible solution to the ever-increasing computational power demand of recent years.
-In the neuromorphic paradigm, computing devices are designed such that computation and memory are colocalized on the physical substrate in the attempt to emulate the brain learning capabilities in materio.
+Neuromorphic computing processors are designed to co-localize computation and memory on the physical substrate and emulate the brain learning capabilities in materia.
+Complete neural networks with dendritic branching have been fabricated with memristive devices and have shown significant power efficiency [24].
 The full potential of biological neural systems is likely achieved because of the interplay between their complex topological structures and the functional dynamics of a very large number of elements evolving on it [78].
-Tree-like structures are currently physically implemented in materio.
-We mention tree structures produced at the critical point of a percolation transition [49, 50] or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [49, 51], but also fractal trees produced in the dielectric breakdown [52], thus, their capabilities could be investigated and enhanced based on this work and future developments.
+Analogously to biological neural systems, structurally efficient routing may play a key role for in-materia computing processors fabricated with self-assembled materials [79, 80].
+Emergent tree-like structures are currently physically fabricated in emerging materials representing potential candidates for neuromorphic computing.
+We mention tree structures produced at the critical point of a percolation transition for randomly dispersed nanoparticles [25, 81], or the electrochemically formed dendrite morphology at the origin of the resistive-switching of some memristive devices [25, 82], and fractal trees produced in the dielectric breakdown [83].
+Thus, their information processing capabilities could be investigated and enhanced based on this work and future developments.
+V Conclusions
+We summarize the contributions included in this work.
+(i) We proposed the Portegys algorithm [27, 28] as an alternative process to produce hierarchies for statistical physics analysis, integrating the functional aspect of a machine-learning algorithm.
+(ii) Multiple phase transitions are identified while the algorithm searches for an efficient internal representation of the dataset.
+(iii) Scale-invariance, i.e.
+power-law Laplacian spectral density [36], is a key feature to constructing efficient tree-networks, capable of combining fast information flow and sufficiently rich internal representation of information.
+Our analysis involved task-agnostic thermodynamic-like efficiency [1], based on general statistical physics arguments, and task-related pattern-matching efficiency, quantifying the trade-off between accuracy and resources used to achieve the task.
+(iv) Studying efficient tree structures mapping the structured disorder of a dataset, we showed that the emergence of complex trees is directly influenced by the complexity of the environmental conditions.
+Namely, the spectral dimension [39, 40, 41], proportional to the exponent of the power-law spectral density, directly encodes the complexity of the disordered environment.
+(v) Therefore, we proposed two novel operational measurements to estimate the dataset complexity: by measuring the spectral dimension of the highly efficient network mapping the dataset; or, equivalently, by measuring the specific heat plateau [34, 35], which reflects the spectral dilation symmetry and the network resolution where information is integrated at a constant rate.
 Conflict of Interest Statement
@@
 Funding
-This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program ”Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No.
+This work was funded by EUs Horizon 2020, from the MSCA-ITN-2019 Innovative Training Networks program “Materials for Neuromorphic Circuits” (MANIC) under the grant agreement No.
 861153, as well as by the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ.
@@
 We are thankful to Jos van Goor for providing some of the scripts that enabled efficient grid search across parameters and to Federico Balducci for the useful discussions.
-Supplemental Data
-The Supplementary Material includes figures referenced in the main text but omitted for clarity.
 Data Availability Statement
-The original contributions presented in the study are included in the article/supplementary material, further inquiries can be directed to the corresponding author/s.
+The original contributions presented in the study are included in the article/supplemental material, further inquiries can be directed to the corresponding author/s.
 Datasets and scripts can be found at https://github.com/CipolliniDavide/Laplacian_Trees.